### v1

# Harnessing Large Language Models for Disaster Management: A Survey

Abstract
Large language models (LLMs) have revolutionized scientific research with their exceptional capabilities and transformed various fields. Among their practical applications, LLMs have been playing a crucial role in mitigating threats to human life, infrastructure, and the environment. Despite growing research in disaster LLMs, there remains a lack of systematic review and in-depth analysis of LLMs for natural disaster management. To address the gap, this paper presents a comprehensive survey of existing LLMs in natural disaster management, along with a taxonomy that categorizes existing works based on disaster phases and application scenarios. By collecting public datasets and identifying key challenges and opportunities, this study aims to guide the professional community in developing advanced LLMs for disaster management to enhance resilience against natural disasters.
Harnessing Large Language Models for Disaster Management: A Survey
Zhenyu Lei   Yushun Dong  Weiyu Li Rong Ding   Qi Wang   Jundong Li University of Virginia, Florida State University, Northeastern University {vjd5zr, jundong}@virginia.edu, yd24f@fsu.edu {weiy.li, ding.ro, q.wang}@northeastern.edu
1 Introduction
Natural disasters have become increasingly frequent and severe, posing unprecedented threats to human life, infrastructure, and the environment Manyena (2006); Yu et al. (2018); Chaudhary and Piracha (2021). The 2010 Haiti earthquake, for instance, led to more than 200,000 fatalities and widespread infrastructure devastation DesRoches et al. (2011). Similarly, the 2020 Australian bushfires have resulted in the deaths of at least 33 people and an estimated one billion animals Deb et al. (2020). The profound impacts of such catastrophic events have highlighted the urgent need for effective disaster management strategies. Recently, large language models (LLMs) have revolutionized the scientific research area due to their unparalleled abilities in contextual understanding, logical generation, and complex problem-solving with textual data and data from various modalities Zhang et al. (2024b, a). These capabilities make LLMs well-suited for natural disaster management, with models capable of analyzing vast disaster real-time data, facilitating dynamic communication with affected communities, and supporting critical decision-making Otal et al. (2024).
Despite the potential of LLMs in disaster management, a systematic review is still missing. This absence hinders researchers and practitioners from identify best practices, uncovering research gaps, and optimizing the deployment of LLMs for effectively addressing disaster-related challenges. To address this gap, this paper systematically surveys the applications of LLMs based on three types of architectures across the four disaster phases: mitigation, preparedness, response, and recovery. We then propose a novel taxonomy that integrates application scenarios, specific tasks, and the architectures of models addressing these tasks. Furthermore, we summarize publicly available datasets, identify key challenges, and investigate avenues for improving the effectiveness, efficiency, and trustworthiness of LLMs. This review aims to inspire and guide AI researchers, policymakers, and practitioners towards developing LLM-aided disaster management frameworks. Our contributions are listed below:
  • We systematically introduce the current explorations handling disaster management with LLMs across four disaster phases.
  • We propose a taxonomy integrating application scenarios, tasks, and model architectures, offering practical and technical insights.
  • We collect key resources such as public datasets and highlight the current challenges and future research opportunities.
Figure 1: Taxonomy of applications of LLMs in disaster management. This survey categorizes the utilization of LLMs across four disaster phases, highlighting specific applications where three types of LLMs (Encoder-based, Decoder-based, and Multimodal LLM) perform tasks such as classification, estimation, extraction, and generation. The chart in the bottom-right corner presents the distribution of surveyed papers across each phase.
2 Background
Disaster management is a multidisciplinary field focused on coordinating resources, expertise, and strategies to address the challenges posed by increasingly severe disasters, minimizing immediate damage while fostering long-term resilience and adaptive recovery. Disaster management consists of four interconnected phases Sun et al. (2020):
  • Mitigation. This phase involves identifying risks and vulnerabilities and implementing proactive measures for disaster prevention.
  • Preparedness. This phase includes comprehensive planning and public education to ensure readiness for upcoming disasters.
  • Response. This phase identifies and addresses immediate needs during a disaster, including rescue and resource distribution.
  • Recovery. This phase involves rebuilding affected areas, addressing both physical and social impacts to facilitate a return to normalcy.
In general, LLMs are able to serve as a general-purpose foundation for building specialized AI tools in disaster management. Here, we categorize LLM architectures into: (1) encoder-based LLM (e.g., BERT Devlin (2018)) primarily for extracting textual features for downstream tasks; (2) decoder-based LLM (e.g., GPT Brown (2020)) for text generation from given textual queries; (3) multimodal LLMs that integrate multiple modalities via alignment Tiong et al. (2022) or feature fusion Madichetty et al. (2021). For disaster management, downstream tasks usually include classification (e.g., disaster detection), estimation (e.g., severity estimation), extraction (e.g., knowledge extraction), and generation (e.g., report generation). Techniques such as fine-tuning, prompting, and additional adapters are commonly employed to adapt LLMs for specific downstream applications.
3 LLM For Disaster Management
Foundation models can be applied to four phases of disaster management, namely mitigation, preparedness, response, and recovery. For each phase, works are organized by application scenarios, tasks, and model architectures. We present the taxonomy in Figure 1 and summaries in Appendix A.
3.1 Disaster Mitigation
Assessing vulnerabilities is a critical aspect of disaster mitigation, where LLMs have shown promising potential. This process typically involves identifying and analyzing infrastructure or communities that are susceptible to disasters.
Vulnerability Classification. Infrastructure Ombudsman has compared the supervised classification with encoder-based LLMs and zero-shot prompt learning with decoder-based LLMs in detecting specific types of concerns about potential infrastructure failures from social media data Chowdhury et al. (2024), facilitating decision-makers in effectively prioritizing resources.
Answer Generation. Apart from the infrastructure vulnerability, decoder-based LLMs can assist in answering community vulnerability-related queries based on retrieval of social vulnerability index (SVI) Martelo and Wang (2024).
3.2 Disaster Preparedness
In the long term, LLMs can play a pivotal role in disaster preparation via (1) public awareness enhancement and consistent (2) disaster forecast. Building on the forecasts, LLMs can support decision-makers in issuing (3) disaster warnings to the public, which can enhance preparedness in the short term. Furthermore, LLMs can assist in well-structured (4) evacuation planning, ensuring safe relocation of individuals and property.
3.2.1 Public Awareness Enhancement
It is critical to enhance public awareness regarding disasters by providing insights and knowledge extracted from previous disaster experiences.
Knowledge Extraction. Encoder-based LLMs are particularly valuable due to their exceptional knowledge extraction capabilities. BERT has been fine-tuned to extract disaster-related knowledge from news and social media Fu et al. (2024), and extensive literature Zhang and Wang (2023). To enhance the logical coherence of extracted entities, Ma et al. propose BERT-BiGRU-CRF for name entity and relation recognition (NER) to construct disaster knowledge graphs Ma et al. (2023). With enhanced reasoning ability, decoder-based LLMs have been utilized, where ChatGLM2 is fine-tuned to extract knowledge triplets from documents to construct a knowledge graph Wu et al. (2024).
Answer Generation. The derived knowledge could be incorporated in decoder-based LLMs’ prompts, which could facilitate answering wildfire Hostetter et al. (2024) and flood risk-related questions Martelo and Wang (2024), and offering personalized mitigation guidance Li et al. (2023). In addition, techniques such as retrieval-augmented generation (RAG) have been utilized for knowledge incorporation, with knowledge retrieved from flood and geographic information system (GIS) knowledge graph Zhu et al. (2024).
3.2.2 Disaster Forecast
Effective disaster preparedness also hinges on accurate and robust forecasting and risk level prediction.
Occurrence Classification. With their extensive internal knowledge, encoder-based LLMs have proven to be powerful tools for classification-based disaster forecasting. For instance, Indra et al. streamline BERT, GRU, and CNN to anticipate disaster Indra and Duraipandian (2023), improving their ability to capture temporal dependencies and local patterns in large-scale social media text. However, textual data is limited due to its subjective and imprecise nature, which leads to multimodal LLMs integrating various data modalities. Zeng et al. combine historical flood data and BERT embeddings extracted from geographical description, then utilize XGBoost to predict the disaster risk Zeng and Bertsimas (2023). Debata et al. further extract features from statistical data with k-nearest neighbors (KNN) for better incorporation. Further advancements incorporate additional satellite imagery features derived from ResNet for more informative disaster prediction Liu and Zhong (2023). To facilitate disaster risk prediction with explicit external knowledge, decoder-based LLMs have been utilized with RAG to retrieve historical flood data from disaster knowledge base to derive risk level and action recommendation Wang et al. (2024).
3.2.3 Disaster Warning
Once a disaster has been foreseen, timely warnings and suggestions are crucial to public safety.
Warning Generation. Decoder-based LLMs have been a valuable tool in generating clear warning messages from rule-based alerts derived from streaming data Chandra et al. (2024), significantly enhancing the responsiveness of warning systems. In addition, LLMs have been utilized to retrieve disaster alerts from National Weather Service’s API to provide detailed, real-time information about upcoming disaster Martelo and Wang (2024).
Image Generation In addition to textual warnings, visual warnings can convey vivid descriptions, reaching a broader audience. To achieve this, multimodal LLMs enhanced by diffusion-based text-to-image generative models can produce vivid visual representations of impending disasters Lubin et al. (2024), improving the clarity of alerts.
3.2.4 Evacuation Planning
Plan Generation
To ensure the safety of individuals and properties, structured evacuation plans have been developed. Leveraging the exceptional reasoning capabilities, decoder-based LLMs have been prompted to plan escapes and provide evacuation suggestions Hostetter et al. (2024), thus enhancing public safety against incoming disasters.
3.3 Disaster Response
With accurate and real-time (1) disaster identification and (2) situation assessment, decision makers can gain essential knowledge to form a response basis. In addition, LLM can facilitate the (3) disaster information coordination, thus boosting a collaborative disaster response among stakeholders. Consequently, decision makers can execute effective actions empowered by LLMs, such as (4) disaster rescue, and offering (5) disaster issue consultation.
3.3.1 Disaster Identification
Effective disaster response begins with accurate and real-time identification to enable efficient interventions Said et al. (2019); Weber et al. (2020), with social media serving as a valuable avenue by providing real-time updates from affected individuals Anderson (2016); Trono et al. (2015).
Relevance Classification. Leveraging their advanced capabilities, LLMs have become prominent disaster detection tools, particularly by performing binary classification for disaster-related posts.
Encoder-based LLMs, with a trainable adapter appended, are widely utilized in this application via fine-tuning Ningsih and Hadiana (2021); Singh et al. (2022). Beyond fine-tuning existing LLMs, researchers have developed novel techniques tailored to specific disaster contexts. CrisisTransformers has pre-trained transformers initialized randomly or with RoBERTa weights on a large dataset with millions of crisis-related tweets before fine-tuning for specific downstream tasks Lamsal et al. (2024a), thus enhancing extensive disaster knowledge incorporation. Additionally, Paul et al. tackle the challenge of limited labeled data by incorporating active learning, which fine-tunes transformer-based representations to prioritize the labeling of the most informative samples Paul et al. (2023).
More innovations focus on framework design, where ensemble models have been employed to integrate the strengths of various LLMs Mukhtiar et al. (2023). Furthermore, additive modifications also enhance LLMs. For instance, several studies append CNN and att-BiLSTM to capture temporal dependencies and local patterns Manthena (2023). In addition, graph neural networks (GNN) can enhance word relation modeling. For instance, Danday et al. construct a word relationship graph from DistilBERT embeddings, which are sent to GNN for classification Danday and Murthy (2022). For better modality fusion, GNom integrates graph features derived from multilingual word graphs and BERT text embeddings with a novel cross-attention mechanism Ghosh et al. (2022).
In addition, advanced decoder-based LLMs such as Llama-2 and GPT-4 demonstrate exceptional contextual understanding ability, being prompted for classification Taghian Dinani et al. (2023).
Image data also offer valuable insights, where multimodal LLMs—combining text and other images—has emerged as a promising approach, enhancing the accuracy of disaster identification. Several approaches combine text features from encoder-based LLMs and image features by aggregation Kamoji et al. (2023); Madichetty et al. (2021). In addition, Koshy et al. append a BiLSTM module after RoBERTa for better sequential modeling Koshy and Elango (2023). To improve modality alignment, Shetty et al. utilize the attention mechanism to fuse text and image features Shetty et al. (2024), while Zhou et al. adopt a Cycle-GAN with mixed fusion strategy to address the multimodal heterogeneity challenges Zhou et al. (2023b). Beyond modality alignment, aligning features across various disaster events is critical in label-scarce scenarios. Jang et al. introduce a semi-supervised minimax entropy domain adaptation framework unifying multimodal features from various disaster events, enabling generalization across diverse contexts Wang and Wang (2022). To combine strengths of diverse LLMs and visual models, the ensemble is further utilized Hanif et al. (2023).
In addition to social media data, other real-time data sources, such as satellite imagery and news articles, can also provide valuable insights. For instance, Jang et al. utilize ResNet to process satellite images and BERT to process descriptions of the disasters from news Jang et al. (2024).
3.3.2 Disaster Situation Assessment
After identifying a disaster, evaluating its severity and spread is crucial for response strategies. This information is often embedded in textual data, such as social media posts reporting situation Saroj and Pal (2020), and news reports offering precise descriptions Houston et al. (2012); Sood et al. (1987).
Situation Classification. Madichetty et al. fine-tune encoder-based LLMs to identify situational posts Madichetty and Sridevi (2021), combining the RoBERTa and SVM representations for binary classification. Raj et al. utilize BERT to filter disaster-related tweets, followed by NER to extract locations, with location counts indicating disaster severity Raj et al. (2023). In addition, visual data from the web and satellites is also integrated and processed by multimodal LLMs to enhance the disaster assessment. Kanth et al. utilize BERT and CNN to filter disaster-related tweets and classify the disaster severity based on the images associated with informative tweets Kanth et al. (2022).
Severity Estimation. Classification only offers a coarse understanding, while severity estimation provides precise quantitative insights. Mousavi et al. utilize chain-of-thoughts (CoT) prompting on decoder-based LLMs (e.g. Gemini) to estimate the earthquake intensity as Modified Mercalli Intensity (MMI) values Mousavi et al. (2024). In addition, multimodal LLMs can be utilized to perform image-related estimation. For instance, FloodDepth-GPT uses prompts to guide GPT-4 to estimate the floodwater depth of the floodwater Akinboyewa et al. (2024) from flood images and descriptions.
Description Generation. Beyond categorical and statistical descriptions, multi-modal LLMs can generate more comprehensible textual situational descriptions. For instance, GPT-4V is prompted to translate situational images to descriptions Hu and Rahnemoonfar (2024), while Wolf et al. prompts GPT-4 to generate assessment descriptions Wolf et al. (2023) with disaster image data.
3.3.3 Disaster Information Coordination
Coordinating disaster-related information is essential for an organized, collaborative response Comfort et al. (2004); Bharosa et al. (2010). Social media plays a pivotal role, with individuals sharing posts on warnings, needs, and other types of information Lindsay (2011); Imran et al. (2015).
Usefulness Estimation. To enhance the accessibility of useful information, Yamamoto et al. leverage encoder-base LLMs (e.g. BERT and RoBERTa) to filter informative tweets by calculating usefulness ratings Yamamoto et al. (2022).
Relevance Classification. However, it requires a pre-defined threshold to determine tweet relevance. Several studies fine-tune encoder-based LLMs for binary relevance classification, which can also be regarded as disaster identification. Furthermore, BERT is applied to classify posts by relevance with multiple relevance levels Blomeier et al. (2024).
Information Classification. To facilitate information dissemination, several studies fine-tuned encoder-based LLMs to classify posts by information types, including actionable types such as "important for managers" Sharma et al. (2021); humanitarian types such as "Injured people" Yuan et al. (2022); and disaster types Liu et al. (2021).
Several studies introduce advanced techniques to improve performance, with some focusing on augmenting data. Boros et al. augment tweets with hashtags and entities, leveraging RoBERTa for classification and tweet prioritizing Boros et al. (2022). Addressing data label scarcity, Li et al. introduce self-training with BerTweet Nguyen et al. (2020) soft-labeling data for training Li et al. (2021).
Novel frameworks have been introduced to improve performance. For instance, Zou et al. utilize a BERT-Bi-LSTM-CNN model to classify tweets into damage types Zou et al. (2024), with better abilities in capturing temporal dependencies and local patterns. In addition, Zahera et al. utilize a graph attention network (GAT) to capture correlations between tweet embeddings and information types Zahera et al. (2021). Furthermore, FF-BERT combines model strengths with an ensemble of BERT and CNN Wilkho et al. (2024).
Other works enhance LLMs by introducing inductive bias. Nguyen et al. leverage BERTweet to extract rationales (classification supporting evidence) from tweets for humanitarian classification Nguyen and Rudra (2022b) and integrate them with attention mechanisms for better fusion Nguyen and Rudra (2023). RACLC Nguyen and Rudra (2022a) further adopts a two-stage framework, using contrastive learning to refine rationale extraction and facilitate classification.
Furthermore, advanced decoder-based LLMs have been utilized, through instruction-tuning Otal and Canbaz (2024); Yin et al. (2024), zero-shot, and few-shot prompting Dinani et al. (2024) for disaster type and humanitarian classification.
Visual data from social media is further incorporated by multimodal LLMs, where several studies have combined features from BERT and visual models via aggregation Zhang et al. (2022); Yu and Wang (2024). Better fusion strategies have been proposed, where Abavisan et al. propose a cross-attention mechanism to fuse DenseNet image features with BERT text embeddings for humanitarian classification Abavisani et al. (2020). Zhou et al. advance this approach by utilizing dual transformers to fuse ALBERT text features and S-CBAM-VGG image features Zhou et al. (2023a). In addition, Basit et al. classify posts into humanitarian or structural types using the "OR" set of text and image classification results Basit et al. (2023).
Need Classification. Social media also allows individuals to express needs, inspiring LLMs for need identification. Encoder-based LLMs have been utilized to detect disaster needs Yang et al. (2024) and rescue request Toraman et al. (2023). Furthermore, MulTMR employs multiple teacher models with task knowledge to fine-tune BERT and RoBERTa for ranking and detecting actionable help requests Vitiugin and Purohit (2024). Responders also share resources on social media. Lamsal et al. have adopted LLMs to match needs with resources through cosine similarity-based retrieval methods where both offer and request posts are embedded by XLM-RoBERTa Conneau (2019), optimizing resource allocation Lamsal et al. (2024b).
Location Extraction. Additionally, efficient processing of posts enhances information dissemination, especially through location extraction. Several studies fine-tune encoder-based LLMs for location reference recognition (LRR), classifying tokens into categories such as "Inside Locations" (ILOC), and "Other Tokens"(O) Mehmood et al. (2024); Suwaileh et al. (2022). Koshy et al. further improve LRR by adding attention and Bi-LSTM layersKoshy and Elango (2024), while others use a BERT-BiLSTM-CRF model to improve the logic consistency of extracted locations Ma et al. (2022); Zhang et al. (2021). External knowledge bases further aid location extraction. For instance, Caillaut et al. utilize cosine similarity to match post entities with knowledge base Caillaut et al. (2024), improving performance. Decoder-based LLMs are also widely applied, where Yu et al. prompt Llama-3 to extract location-relevant words Yu and Wang (2024). External knowledge has been utilized for more accurate location extraction. Hu et al. leverage GPT-4 for detailed location descriptions, incorporating geo-knowledge within the prompts Hu et al. (2023). In addition, GeoChatGPT augments LLMs with geographic data and Object Character Recognition (OCR)-based object descriptions to predict disaster locations Firmansyah et al. (2024).
Summary Extraction. Furthermore, extracting and summarizing critical posts provides a macro understanding during disasters. Nguyen et al utilize integer linear programming (ILP) with encoder-based LLMs to identify critical posts and rationales for summaries Nguyen and Rudra (2022a); Nguyen et al. (2022). Garg et al. calculate salience scores using BERT embeddings and word2vec key-phrase embeddings from rapid automatic keyword extraction (RAKE)Garg et al. (2024). Decoder-based LLMs go further by generating summaries from retrieved tweets. For instance, Vitiugin et al. rank important tweets utilizing LSTM and then use a T5 model to generate the summaries based on top retrieved tweets Vitiugin and Castillo (2022).
Report Generation. Organizations also use decoder-based LLMs for structured disaster reporting, especially through zero-shot search-based generation Colverd et al. (2023). In addition, Seeberger et al. utilize multiple LLMs each for document retrieval, reranking, and instruction-following summarization Seeberger and Riedhammer (2024b), while chain-of-thought (CoT) reasoning is further utilized Pereira et al. (2023). Building on these methods, Crisis2Sum further organizes query-related information into topic clusters summarized by Mixtral Jiang et al. (2024), and produces a final report by combining key nuggets from each cluster Seeberger and Riedhammer (2024a).
3.3.4 Disaster Rescuing
Grounded in a comprehensive disaster situation understanding, disaster rescuing aims to save lives and property through timely, coordinated actions.
Plan Generation. Effective rescue operations typically require comprehensive rescue plans. Grounded with rich disaster-related knowledge, decoder-based LLMs have been prompted to generate actionable response plans that provide essential guidance Goecks and Waytowich (2023).
Code Generation. Once a plan is established, decoder-based LLMs can assist in execution performed by organizations and rescue teams, such as assisting in guiding robotic systems by translating verbal inputs into actionable operational commands using RAG Panagopoulos et al. (2024).
3.3.5 Disaster Issue Consultation
During disasters, affected individuals and organizations often seek reliable guidance. Disaster issue consultation offers advice, safety updates, and expert guidance, thus helping them access resources, evaluate options, and make decisions Jiang (2024).
Answer Generation. Decoder-based LLMs have been prompted to generate answers for frequently asked questions and offer guidance Rawat (2024); Chen and Fang (2024). To mitigate hallucination, LLMs integrate RAG with verified disaster-related documents. For instance, WildfireGPT retrieves wildfire literature and data to augment the prompts Xie et al. (2024). Chen et al. guide LLM reasoning over KG by a prompt chain, adding structured knowledge Chen et al. (2024). Unlike traditional RAG approaches that rely on prompt learning, Xia et al. integrate fine-tuning with RAG to enhance LLMs by combining implicit and external knowledge. Xia et al. (2024). Additionally, multi-modal LLMs can integrate textual and visual data. For instance, several VQA models (e.g., Plug-and-Play VQA) Tiong et al. (2022) have been prompted for zero-shot visual question answering for disaster scenarios Sun et al. (2023). To interpret complex user queries, ADI further introduces sequential modular tools including VLMs, object detection models, and semantic segmentation models Liu et al. (2024). In addition, FloodLense integrates ChatGPT and diffusion models to highlight disaster-affected areas from images to improve flood-related geographical question answering Kumbam and Vejre (2024).
3.4 Disaster Recovery
LLMs can perform accurate (1) disaster impact assessment, a critical step for disaster recovery. With a comprehensive impact understanding, LLMs can assist decision makers in (2) recovery plans generation. In addition, disaster responders have utilized LLMs for (3) recovery process tracking consistently, ensuring the recovery effectiveness.
3.4.1 Disaster Impact Assessment
Accurate assessment of the damage degree across physical and social dimensions helps prioritize recovery efforts effectively. Physically, disasters disrupt infrastructure, damage property, and threaten human lives, requiring robust frameworks for the identification and estimation of damage and loss.
Damage Classification. Various encoder-based LLMs have been employed to identify and categorize disaster damage. For instance, Malik et al. utilize BERT to identify the damage-related posts and categorize them as "human damage" or "infrastructure damage" Malik et al. (2024). Chen et al. further categorize damage into four types such as "Water or Power Supply" with BERT Chen and Lim (2021). In addition, Zou et al. propose a BERT-BiLSTM-Sit-CNN framework for damage-related posts identification and damage-type classification with improved textual understanding Zou et al. (2024). Furthermore, several studies utilize LLMs to measure damage severity, where Jeba et al. leverage BERT for damage impact degree classification on social media posts and news  Jeba et al. (2024).
Damage Estimation. Damage severity could be better modeled quantitatively by estimation with a fine-grained understanding. Chen et al. calculate the damage severity scores based on the similarities between post tokens and several seed words’ embeddings, where both embeddings are derived from encoder-based LLMs Chen and Lim (2021).
Answer Generation. In addition, decoder-based LLMs can answer specific assessment questions. For instance, Ziaullah et al. prompt LLMs with RAG to obtain the operation status of critical infrastructure facilities using social media data Ziaullah et al. (2024). Multimodal LLMs further incorporate remote sensing data, with Estevao et al. prompting GPT-4o to generate damage assessment using building images Estêvão (2024). To improve modality alignment, SAM-VQA employs a supervised attention-based VLM to integrate image and question features for VQA tasks Sarkar et al. (2023). Additionally, axillary tasks have been utilized to improve the VQA performance, where DATWEP dynamically balances the importance of segmentation and VQA tasks, adjusting class weights during training Alsan and Arsan (2023).
Statistic Extraction. Furthermore, decoder-based LLMs are prompted to extract fatalities Hou and Xu (2022) from social media via few-shot learning, thus offering timely insights into human loss.
Sentiment Classification. From the social dimension, disasters influence public sentiment and behavior, where encoder-based LLMs Han et al. (2024a); Berbère et al. (2023) have been fine-tuned to classify social media posts into positive and negative emotions. In addition, Li et al. utilize decoder-based LLM (e.g. GPT 3.5) to classify the posts into five emotional types such as "panic" and "sadness" with zero-shot prompting Li et al. (2025), thus aiding responders in understanding and addressing the disaster’s emotional impact.
3.4.2 Recovery Plan Generation
Based on impact assessment, a recovery plan is developed to rebuild infrastructure, restore services, and enhance resilience Hallegatte et al. (2018).
Plan Generation. Decoder-based LLMs have been utilized in certain recovery scenarios by generating recovery and reconstruction plans. For instance, ChatGPT has been prompted to build disaster recovery plans for business restoration White and Liptak (2024); Lakhera (2024).
3.4.3 Recovery Process Tracking
Continuous tracking of the recovery process ensures that progress aligns with the planned timeline and goals and decision-makers can adapt recovery strategies to meet evolving needs.
Sentiment Classification. Encoder-based LLM (e.g. BERTweet) has been utilized to assess public sentiment throughout the post-disaster recovery period CONTRERAS et al. , thus enabling responders to tailor their recovery efforts to address emotional needs of affected population effectively.
4 Datasets
Numerous disaster-related datasets have been utilized to evaluate LLMs in disaster management, and we present a comprehensive list of publicly available datasets in Appendix B.
Classification datasets are primarily designed with textual inputs from platforms like Twitter and news outlets, categorizing data by informativeness (relevance) Olteanu et al. (2014), humanitarian types Imran et al. (2016), damage levels Alam et al. (2021b), and other attributes. Some also include visual data, such as satellite imagery or social media images Alam et al. (2018). Models are evaluated with metrics like accuracy and F1 score.
Estimation datasets include textual and visual inputs and provide quantitative labels such as flood depths Akinboyewa et al. (2024), evaluated using metrics such as Mean Absolute Error (MAE).
Generation datasets are also extensively used, primarily in two categories: question answering and summarization. Question-answering datasets provide disaster-related questions paired with crowd-source annotated answers Rawat (2024). Multimodal question-answering datasets, which incorporate disaster-related images as contextual information, are widely utilized as well Sun et al. (2023). For summarization task, large sets of documents are used as inputs, with reference summaries curated by domain experts McCreadie and Buntain (2023). Both types of generation tasks are evaluated using metrics like ROUGE and BLEU.
Extraction datasets identify and label specific elements within sentence, such as keywords Nguyen and Rudra (2022a) and locations Suwaileh et al. (2022), labeling tokens as “outside,” “start,” or “end” to specify their extraction status. These datasets are primarily used for token-level classification tasks, and is measured with classification metrics.
5 Challenges and Opportunities
Large language models (LLMs) show promise in disaster management but face key limitations. Applications are disproportionately focused on disaster response, as illustrated in Figure 1, highlighting the need to expand their use across all disaster phases. Additionally, there is a gap between researchers focusing on novel architectures and practitioners needing disaster-specific solutions, necessitating LLMs tailored to various challenges.
Dataset Scarcity. Current datasets are heavily skewed toward classification tasks, leaving other critical areas underrepresented. Creating high-quality datasets requires significant effort and resources, making the process both labor- and data-intensive. Innovative solutions like synthetic data generation Kalluri et al. (2024) are promising to expand coverage across disaster-related scenarios.
Efficient Deployment. large-scale LLMs face latency and computational challenges, hindering real-time decision-making in emergent disaster scenarios. Lightweight models provide a faster, resource-efficient alternative Saleem et al. (2024), but they often sacrifice accuracy and robustness. Advancing these models to balance efficiency with reliability is crucial for effective disaster management.
Robust Generation. Decoder-based LLMs may hallucinate, where outputs can be factually inaccurate, posing serious risks in disaster contexts, such as false evacuation routes, resource misallocation, and potential loss of lives. Integrating RAG with external knowledge bases Colverd et al. (2023), extensive domain-specific training Lamsal et al. (2024a), and uncertainty estimation Xu et al. can potentially deal with hallucinated outputs.
6 Conclusion
This paper surveys LLMs in disaster management across four disaster phases, with a taxonomy integrating application scenarios, specific tasks, and the architectures of models addressing these tasks. With introducing publicly available datasets and identifying challenges, we inspire the collaborative efforts between AI researchers and decision-makers, in order to fully leveraging the power of LLMs for more resilient communities and proactive disaster management practices.
Limitations
This work focuses exclusively on disaster management applications only where existing LLMs have been utilized, leaving out other potential scenarios that have not yet been explored by current LLM research. While these unexplored areas hold significant promise for future advancements, they are beyond the scope of this work due to space constraints. Additionally, we include only a subset of datasets employed in existing studies, prioritizing those that are easily accessible. Many datasets are either not open-sourced, have restrictive access policies, or lack assured quality, making them less viable for reproducibility and further research.
Appendix A Summary of Papers
A.1 Summary Table
Table LABEL:tab:model provides a summary of the surveyed papers with their disaster phases, application scenarios, specific tasks, and architecture types.
A.2 Pipeline Illustration
In this section, we present Figure 2 to illustrate the use of LLMs in disaster management. The figure outlines the major pipelines of three LLM architectures—encoder-based, decoder-based, and multimodal—applied across four task types in our survey: classification, extraction, estimation, and generation, offering critical insights into their mechanisms for disaster management.
A.3 Statistics
To provide a comprehensive overview of the current state of LLMs in disaster management, we present statistics from the surveyed papers, highlighting a significant gap between the NLP and disaster management communities. This gap underscores the pressing need for better interdisciplinary collaboration to bridge the two fields and realize the potential of LLMs in addressing disaster-related challenges.
Figure 3 depicts the counts of publications leveraging existing LLMs and developing new frameworks, which indicates that most studies are heavily application-focused, relying on fine-tuning or prompting existing LLMs for disaster management tasks. In addition, while some efforts have yielded valuable insights, the majority of these studies are predominantly concentrated on the response phase of disaster management, with limited exploration across other disaster management scenarios. Figure 4 illustrates the publication distribution across academic venues, revealing that relatively few disaster management papers are published in NLP or AI-specific conferences or journals. These findings reflect limited engagement from the LLM research community in this domain, highlighting a need to increase awareness and focus within the field.
Appendix B Datasets
Table LABEL:tab:dataset provides a summary of existing publicly available datasets. For classification tasks, we exclude datasets limited to a single disaster type that are already incorporated into comprehensive benchmarks such as CrisisBench Alam et al. (2021b).
B.1 Classification Datasets
  • CrisisLexT6 Olteanu et al. (2014): This dataset is designed for relevance classification. It contains data from six crisis events between October 2012 and July 2013.
  • CrisisLexT26 Olteanu et al. (2015): This dataset is an updated version of CrisisLexT6, which contains public data from 26 crisis events in 2012 and 2013 with relevance information and six humanitarian categories.
  • CrisisNLP Imran et al. (2016): This dataset is a large-scale dataset including classes from humanitarian disaster responses and classes related to health emergencies. It is collected from 19 different disaster events that happened between 2013 and 2015.
  • SWDM2013 Imran et al. (2013): This dataset is utilized for relevance classification consisting of tweets from two events: (i) the Joplin collection contains tweets from the tornado that struck Joplin, Missouri on May 22, 2011; (ii) The Sandy collection contains tweets collected from Hurricane Sandy that hit Northeastern US on Oct 29, 2012.
  • ISCRAM2013 Imran et al. (2013): This dataset consists of tweets collected from the same events as in SWDM2013, containing both relevance and humanitarian categories.
  • Disaster Response Data (DRD) Alam et al. (2021b): This dataset consists of tweets collected during various crisis events that took place in 2010 and 2012. This dataset is annotated using 36 classes that include relevance as well as humanitarian categories.
  • Disasters on Social Media (DSM) Alam et al. (2021b): This dataset comprises 10K tweets annotated with relevance labels.
  • AIDR Imran et al. (2014): This dataset contains data obtained from the AIDR system on September 25, 2013, collecting tweets using hashtags such as "#earthquake". It is utilized for relevance and humanitarian classification.
  • CrisisMMD Alam et al. (2018): This dataset is a multimodal and multitask dataset comprising $16k$ labeled tweets and corresponding images. Tweets have been sourced from seven natural disaster events that took place in 2017. Each sample is annotated with relevance, humanitarian (eight classes), and damage severity categories (mild, severe, and none).
  • Multi-Crisis Sánchez et al. (2023): This dataset was proposed to evaluate transfer learning scenarios where data from high-resource languages (e.g., English) is used to classify messages in low-resource languages (e.g., Spanish, Italian) and unseen crisis domains, with relevance and humanitarian categories. It is collected from 7 existing datasets, 53 crisis events, and contains 9 domains.
  • CrisisBench Alam et al. (2021b): This dataset is a comprehensive benchmark consolidated from 9 existing datasets, utilized for relevance and humanitarian classification.
  • Eyewitness Messages Zahra et al. (2020): This dataset is proposed to identity disaster eyewitness related tweets and classifies them into three types: direct eyewitnesses, indirect eyewitnesses, and vulnerable eyewitnesses where people anticipate a disaster and were present in the region for which disaster warnings were issued. It is comprised of $14k$ data from earthquakes, hurricanes, and wildfires.
  • TREC Incident Streams McCreadie et al. (2019): This dataset has been developed as part of the TREC-IS 2018 evaluation challenge and consists of $20k$ tweets labeled for actionable information identification and information criticality assessment.
  • HumAID Alam et al. (2021a): This dataset contains $77k$ labeled tweets, which are sampled from 24 million tweets collected during 19 disasters between 2016 and 2019, including hurricanes, earthquakes, wildfires, and floods. It is balanced in terms of disaster types and contains 7 humanitarian categories.
  • EPIC: This dataset contains data primarily collected from Hurricane Sandy, including tweets from 93 users across four annotation schemes, with data spanning three weeks around the hurricane’s landfall. It is used for relevance and humanitarian classification.
Figure 3: Counts of publications leveraging existing LLMs and developing new models across four phases.
B.2 Estimation Datasets
  • Did You Feel It (DYFI) Atkinson and Wald (2007): This dataset contains ground shaking intensity and geographic distribution information, collected from post-earthquake reports (through $750k$ online questionnaire responses) from individuals who experienced the event.
  • FloodDepth Akinboyewa et al. (2024): This dataset contains 150 flood photos collected online, which is utilized to estimate floodwater depth based on different reference objects, including stop signs, vehicles, and humans.
  • Behavioral Facilitation (BF) Yamamoto et al. (2022): This datasets, collected after the 2018 Hokkaido earthquake, contains data labeled with usefulness ratings with respect to behavioral facilitation information.
Figure 4: Publication distribution across various academic venues, with a focus on model design on the left and application-based research on the right.
B.3 Extraction Datasets
  • Fu et al. (2024): This dataset contains county-level data from news media collected during urban flood events from 2000 to 2022. It is utilized to extract information such as the time and location of disasters.
  • Ma et al. (2023): This dataset is used for entity and relation extraction, containing $5,560$ annotated data, $12,980$ entities, and $6,895$ relations, from report during geological hazards.
  • DisasterMM Andreadis et al. (2022): This dataset is collected from Twitter by searching for flood-related keywords. It comprises two subsets: RCTP, containing 6,672 tweets for relevance classification; LETT, consisting of 4,992 tweets used for location extraction, where words are labeled with “B-LOC” for the first word of a sequence that refers to a location, “I-LOC” for the subsequent word of a sequence that refers to a location, and “O” for any non-location word.
  • Suwaileh et al. (2022): This dataset contain $22k$ crisis-related tweets from several disasters including floods, earthquakes, and hurricane. It is labeled with location-related tags such as "inLOC" and "outLOC".
  • Re´SoCIO Caillaut et al. (2024): This dataset is constructed by merging Wikipedia datasets and multiple disaster-related datasets, annotated with a set of 9 NER labels with different types of information.
  • Nguyen and Rudra (2022a): This dataset contains tweet data with annotated rationales from 4 subsets of CrisisNLP. It is utilized for rationale extraction, and the extracted rationales can assist in disaster classification.
B.4 Generation Datasets
  • Vitiugin and Castillo (2022): This dataset is utilized to generate summaries for various disaster events, where the ground truth is the official report of each event.
  • CrisisFACTS McCreadie and Buntain (2023): This dataset is a multiple-stream dataset with data related to 8 crisis events from various platforms. It is utilized to consume daily multi-platform streams and produce summaries for a given information need, such as "Have airports closed?".
  • DisasterQA Rawat (2024): This dataset includes disaster-related multiple choice questions from 7 different sources, examples could be "What causes a tsunami?".
  • FFD-IQA Sun et al. (2023): This dataset includes 2,058 images and 22,422 question-meta ground truth pairs related to the safety of individuals trapped in disaster sites and the availability of emergency services. In total, three types of questions are included: free-form questions, multiple choice questions, and yes-no questions.
  • FloodNet Rahnemoonfar et al. (2021): The dataset contains 4,500 question-image pairs collected after Hurricane Harvey. Questions are related to the building, road, and the entire image, and are divided into 4 groups: "Simple Counting", "Complex Counting", "yes/no", and "Condition Recognition".
Table 1: Summary of LLMs in disaster management with their disaster phases, application scenarios, specific tasks, and architecture types. "Arch": Type of LLM architectures used; "NM": Whether the paper presents novel methods.
Paper Phase Application Task Arch NM
Chowdhury et al. (2024) Mitigation Vulnerability Assessment Vulnerability Classification Decoder No
Martelo and Wang (2024) Mitigation Vulnerability Assessment Answer Generation Decoder Yes
Fu et al. (2024) Preparedness Public Awareness Enhancement Knowledge Extraction Encoder No
Zhang and Wang (2023) Preparedness Public Awareness Enhancement Knowledge Extraction Encoder No
Ma et al. (2023) Preparedness Public Awareness Enhancement Knowledge Extraction Encoder Yes
Wu et al. (2024) Preparedness Public Awareness Enhancement Knowledge Extraction Decoder No
Hostetter et al. (2024) Preparedness Public Awareness Enhancement Answer Generation Decoder No
Martelo and Wang (2024) Preparedness Public Awareness Enhancement Answer Generation Decoder No
Li et al. (2023) Preparedness Public Awareness Enhancement Answer Generation Decoder No
Indra and Duraipandian (2023) Preparedness Disaster Forecast Occurrence Classification Encoder Yes
Zeng and Bertsimas (2023) Preparedness Disaster Forecast Occurrence Classification Multimodal Yes
Liu and Zhong (2023) Preparedness Disaster Forecast Occurrence Classification Multimodal Yes
Wang et al. (2024) Preparedness Disaster Forecast Occurrence Classification Decoder No
Chandra et al. (2024) Preparedness Disaster Warning Warning Generation Decoder No
Martelo and Wang (2024) Preparedness Disaster Warning Warning Generation Decoder No
Lubin et al. (2024) Preparedness Disaster Warning Image Generation Multimodal Yes
Hostetter et al. (2024) Preparedness Evacuation Planning Plan Generation Decoder No
Ningsih and Hadiana (2021) Response Disaster Identification Relevance Classification Encoder No
Madichetty and Madisetty (2023) Response Disaster Identification Relevance Classification Encoder No
Singh et al. (2022) Response Disaster Identification Relevance Classification Encoder No
Ranade et al. (2021) Response Disaster Identification Relevance Classification Encoder No
Powers et al. (2023) Response Disaster Identification Relevance Classification Encoder No
Jain et al. (2019) Response Disaster Identification Relevance Classification Encoder No
Duraisamy and Natarajan (2024) Response Disaster Identification Relevance Classification Encoder No
Ullah et al. (2023) Response Disaster Identification Relevance Classification Encoder No
Li and Chen (2024) Response Disaster Identification Relevance Classification Encoder No
Zhao et al. (2024) Response Disaster Identification Relevance Classification Encoder No
Karanjit et al. (2024) Response Disaster Identification Relevance Classification Encoder No
Pabari et al. (2023) Response Disaster Identification Relevance Classification Encoder No
de Bruijn et al. (2019) Response Disaster Identification Relevance Classification Encoder No
Zhao et al. (2024) Response Disaster Identification Relevance Classification Encoder No
Wang et al. (2021) Response Disaster Identification Relevance Classification Encoder No
Habib et al. (2024) Response Disaster Identification Relevance Classification Encoder No
Liu et al. (2021) Response Disaster Identification Relevance Classification Encoder No
Fontalis et al. (2023) Response Disaster Identification Relevance Classification Encoder No
Mehmood et al. (2024) Response Disaster Identification Relevance Classification Encoder No
Paul et al. (2023) Response Disaster Identification Relevance Classification Encoder Yes
Lamsal et al. (2024a) Response Disaster Identification Relevance Classification Encoder Yes
Manthena (2023) Response Disaster Identification Relevance Classification Encoder Yes
Danday and Murthy (2022) Response Disaster Identification Relevance Classification Encoder Yes
Ghosh et al. (2022) Response Disaster Identification Relevance Classification Encoder Yes
Taghian Dinani et al. (2023) Response Disaster Identification Relevance Classification Decoder No
Kamoji et al. (2023) Response Disaster Identification Relevance Classification Multimodal Yes
Madichetty et al. (2021) Response Disaster Identification Relevance Classification Multimodal Yes
Koshy and Elango (2023) Response Disaster Identification Relevance Classification Multimodal Yes
Shetty et al. (2024) Response Disaster Identification Relevance Classification Multimodal Yes
Zhou et al. (2023b) Response Disaster Identification Relevance Classification Multimodal Yes
Yu and Wang (2024) Response Disaster Identification Relevance Classification Multimodal Yes
Zhang et al. (2022) Response Disaster Identification Relevance Classification Multimodal Yes
Kota et al. (2022) Response Disaster Identification Relevance Classification Multimodal Yes
Wang and Wang (2022) Response Disaster Identification Relevance Classification Multimodal Yes
Hanif et al. (2023) Response Disaster Identification Relevance Classification Multimodal Yes
Jang et al. (2024) Response Disaster Identification Relevance Classification Multimodal Yes
Madichetty and Sridevi (2021) Response Disaster Situation Assessment Situation Classification Encoder Yes
Raj et al. (2023) Response Disaster Situation Assessment Situation Classification Encoder Yes
Kanth et al. (2022) Response Disaster Situation Assessment Situation Classification Multimodal Yes
Mousavi et al. (2024) Response Disaster Situation Assessment Severity Estimation Decoder No
Akinboyewa et al. (2024) Response Disaster Situation Assessment Severity Estimation Multimodal No
Hu and Rahnemoonfar (2024) Response Disaster Situation Assessment Description Generation Multimodal No
Wolf et al. (2023) Response Disaster Situation Assessment Description Generation Multimodal No
Yamamoto et al. (2022) Response Disaster Information Coordination Usefulness Estimation Encoder No
Blomeier et al. (2024) Response Disaster Information Coordination Relevance Classification Encoder No
Adesokan et al. (2023) Response Disaster Information Coordination Information Classification Encoder No
Wahid et al. (2022) Response Disaster Information Coordination Information Classification Encoder No
Chandrakala and Raj (2022) Response Disaster Information Coordination Information Classification Encoder No
Naaz et al. (2021) Response Disaster Information Coordination Information Classification Encoder No
Du et al. (2023) Response Disaster Information Coordination Information Classification Encoder No
Adesokan et al. (2023) Response Disaster Information Coordination Information Classification Encoder No
Han et al. (2024b) Response Disaster Information Coordination Information Classification Encoder No
Sharma et al. (2021) Response Disaster Information Coordination Information Classification Encoder No
Yuan et al. (2022) Response Disaster Information Coordination Information Classification Encoder No
Liu et al. (2021) Response Disaster Information Coordination Information Classification Encoder No
Boros et al. (2022) Response Disaster Information Coordination Information Classification Encoder Yes
Li et al. (2021) Response Disaster Information Coordination Information Classification Encoder Yes
Zou et al. (2024) Response Disaster Information Coordination Information Classification Encoder Yes
Zahera et al. (2021) Response Disaster Information Coordination Information Classification Encoder Yes
Wilkho et al. (2024) Response Disaster Information Coordination Information Classification Encoder Yes
Nguyen and Rudra (2022b) Response Disaster Information Coordination Information Classification Encoder Yes
Nguyen and Rudra (2023) Response Disaster Information Coordination Information Classification Encoder Yes
Nguyen and Rudra (2022a) Response Disaster Information Coordination Information Classification Encoder Yes
Otal and Canbaz (2024) Response Disaster Information Coordination Information Classification Decoder No
Yin et al. (2024) Response Disaster Information Coordination Information Classification Decoder No
Dinani et al. (2024) Response Disaster Information Coordination Information Classification Decoder No
Zhang et al. (2022) Response Disaster Information Coordination Information Classification Multimodal Yes
Yu and Wang (2024) Response Disaster Information Coordination Information Classification Multimodal Yes
Shetty et al. (2024) Response Disaster Information Coordination Information Classification Multimodal Yes
Abavisani et al. (2020) Response Disaster Information Coordination Information Classification Multimodal Yes
Zhou et al. (2023a) Response Disaster Information Coordination Information Classification Multimodal Yes
Basit et al. (2023) Response Disaster Information Coordination Information Classification Multimodal Yes
Yang et al. (2024) Response Disaster Information Coordination Need Classification Encoder No
Toraman et al. (2023) Response Disaster Information Coordination Need Classification Encoder No
Zhou et al. (2022) Response Disaster Information Coordination Need Classification Encoder No
Vitiugin and Purohit (2024) Response Disaster Information Coordination Need Classification Encoder Yes
Conneau (2019) Response Disaster Information Coordination Need Classification Encoder Yes
Lamsal et al. (2024b) Response Disaster Information Coordination Need Classification Encoder Yes
Mehmood et al. (2024) Response Disaster Information Coordination Location Extraction Encoder No
Suwaileh et al. (2022) Response Disaster Information Coordination Location Extraction Encoder No
Koshy and Elango (2024) Response Disaster Information Coordination Location Extraction Encoder Yes
Ma et al. (2022) Response Disaster Information Coordination Location Extraction Encoder Yes
Zhang et al. (2021) Response Disaster Information Coordination Location Extraction Encoder Yes
Caillaut et al. (2024) Response Disaster Information Coordination Location Extraction Encoder Yes
Yu and Wang (2024) Response Disaster Information Coordination Location Extraction Decoder No
Hu et al. (2023) Response Disaster Information Coordination Location Extraction Decoder No
Firmansyah et al. (2024) Response Disaster Information Coordination Location Extraction Decoder No
Nguyen and Rudra (2022a) Response Disaster Information Coordination Summary Extraction Encoder Yes
Nguyen et al. (2022) Response Disaster Information Coordination Summary Extraction Encoder Yes
Garg et al. (2024) Response Disaster Information Coordination Summary Extraction Encoder Yes
Vitiugin and Castillo (2022) Response Disaster Information Coordination Summary Extraction Decoder Yes
Colverd et al. (2023) Response Disaster Information Coordination Report Generation Decoder No
Pereira et al. (2023) Response Disaster Information Coordination Report Generation Decoder No
Seeberger and Riedhammer (2024b) Response Disaster Information Coordination Report Generation Decoder Yes
Seeberger and Riedhammer (2024a) Response Disaster Information Coordination Report Generation Decoder Yes
Goecks and Waytowich (2023) Response Disaster Rescuing Plan Generation Decoder No
Panagopoulos et al. (2024) Response Disaster Rescuing Code Generation Decoder No
Rawat (2024) Response Disaster Issue Consultation Answer Generation Decoder No
Chen and Fang (2024) Response Disaster Issue Consultation Answer Generation Decoder No
Xie et al. (2024) Response Disaster Issue Consultation Answer Generation Decoder No
Chen et al. (2024) Response Disaster Issue Consultation Answer Generation Decoder Yes
Xia et al. (2024) Response Disaster Issue Consultation Answer Generation Decoder Yes
Sun et al. (2023) Response Disaster Issue Consultation Answer Generation Multimodal No
Liu et al. (2024) Response Disaster Issue Consultation Answer Generation Multimodal Yes
Kumbam and Vejre (2024) Response Disaster Issue Consultation Answer Generation Multimodal Yes
Malik et al. (2024) Recovery Disaster Impact Assessment Damage Classification Encoder No
Chen and Lim (2021) Recovery Disaster Impact Assessment Damage Classification Encoder No
Jeba et al. (2024) Recovery Disaster Impact Assessment Damage Classification Encoder No
Zou et al. (2024) Recovery Disaster Impact Assessment Damage Classification Encoder Yes
Chen and Lim (2021) Recovery Disaster Impact Assessment Damage Estimation Encoder Yes
Ziaullah et al. (2024) Recovery Disaster Impact Assessment Answer Generation Decoder No
Estêvão (2024) Recovery Disaster Impact Assessment Answer Generation Multimodal No
Sarkar et al. (2023) Recovery Disaster Impact Assessment Answer Generation Multimodal No
Alsan and Arsan (2023) Recovery Disaster Impact Assessment Answer Generation Multimodal No
Hou and Xu (2022) Recovery Disaster Impact Assessment Statistic Extraction Decoder No
Han et al. (2024a) Recovery Disaster Impact Assessment Sentiment Classification Encoder No
Alharm and Naim Recovery Disaster Impact Assessment Sentiment Classification Encoder No
Zhang and Ma (2023) Recovery Disaster Impact Assessment Sentiment Classification Encoder No
Varghese et al. (2024) Recovery Disaster Impact Assessment Sentiment Classification Encoder No
Berbère et al. (2023) Recovery Disaster Impact Assessment Sentiment Classification Encoder No
Li et al. (2025) Recovery Disaster Impact Assessment Sentiment Classification Decoder No
White and Liptak (2024) Recovery Recovery Plan Generation Plan Generation Decoder No
Lakhera (2024) Recovery Recovery Plan Generation Plan Generation Decoder No
CONTRERAS et al. Recovery Recovery Process Tracking Sentiment Classification Encoder No
Table 2: Summary of publicly available datasets utilized in disaster management. For Application, "DI": Disaster Identification; "DInf": Disaster Information Coordination; "DIC": Disaster Issue Consultation; "DSA": Disaster Situation Assessment; "PAE": Public Awareness Enhancement; "DIA": Disaster Impact Assessment. For Disaster Type, "Mix" denotes the datasets contain various types of disasters.
Dataset Phase Application Task Disaster Type Modality Used in #Sample
CrisisLexT6 Olteanu et al. (2014) Response DI Classification Mix Text McDaniel et al. (2024) 60,082
CrisisLexT26 Olteanu et al. (2015) Response DI, DInf Classification Mix Text McDaniel et al. (2024) 27,933
CrisisNLP Imran et al. (2016) Response DI, DInf Classification Mix Text Taghian Dinani et al. (2023) 52,656
SWDM13 Imran et al. (2013) Response DI, DInf Classification Mix Text McDaniel et al. (2024) 1,543
ISCRAM2013 Imran et al. (2013) Response DI, DInf Classification Mix Text McDaniel et al. (2024) 3,617
DRD Alam et al. (2021b) Response DI, DInf Classification Mix Text McDaniel et al. (2024) 26,235
DSM Alam et al. (2021b) Response DI Classification Mix Text McDaniel et al. (2024) 10,876
AIDR Imran et al. (2014) Response DI, DInf Classification Mix Text McDaniel et al. (2024) 7,411
CrisisMMD Alam et al. (2018) Response DI, DInf Classification Mix Text, Image Jain et al. (2024) 16,058
Multi-Crisis Sánchez et al. (2023) Response DI, DInf Classification Mix Text Sánchez et al. (2023) 164,625
CrisisBench Alam et al. (2021b) Response DI, DInf Classification Mix Text McDaniel et al. (2024) 109,796
Eyewitness Messages Zahra et al. (2020) Response DInf Classification Mix Text Zahra et al. (2020) 14,000
TREC Incident Streams McCreadie et al. (2019) Response DI, DInf Classification Mix Text Khattar and Quadri (2022) 19,784
HumAID Alam et al. (2021a) Response DInf Classification Mix Text Basit et al. (2023) 77,000
EPIC Stowe et al. (2018) Response DI, DInf Classification Mix Text Adesokan et al. (2023) 3469
Did You Feel It (DYFI) Mousavi et al. (2024) Response DSA Estimation Earthquake Text Mousavi et al. (2024) 750,000
FloodDepth Akinboyewa et al. (2024) Response DSA Estimation Flood Text, Image Akinboyewa et al. (2024) 150
Behavioral Facilitation (BF) Yamamoto et al. (2022) Response DInf Estimation Earthquake Text Yamamoto et al. (2022) 1,400
Fu et al. (2024) Preparedness PAE Extraction Flood Text Fu et al. (2024) 633
Ma et al. (2023) Preparedness PAE Extraction Landslide Text Ma et al. (2023) 5,560
DisasterMM Andreadis et al. (2022) Response DI, DInf Classification, Extraction Flood Text Mehmood et al. (2024) 6,672, 4,992
Suwaileh et al. (2022) Response DInf Extraction Mix Text Suwaileh et al. (2022) 22,137
Re´SoCIO Caillaut et al. (2024) Response DInf Extraction Flood Text Caillaut et al. (2024) 4,617
 Nguyen and Rudra (2022a) Response DInf Extraction Mix Text Nguyen and Rudra (2022a) 32
 Vitiugin and Castillo (2022) Response DInf Generation Mix Text Vitiugin and Castillo (2022) 5,791
CrisisFACTS McCreadie and Buntain (2023) Response DIC Generation Mix Text Pereira et al. (2023) 748,466
DisasterQA Rawat (2024) Response PAE, DIC Generation Mix Text Rawat (2024) 707
FFD-IQA Sun et al. (2023) Response DIC Generation Flood Text, Image Sun et al. (2023) 22,422
FloodNet Rahnemoonfar et al. (2021) Recovery DIA Generation Flood Text, Image Sarkar et al. (2023) 4,500

### v2

@@
 Abstract
-Large language models (LLMs) have revolutionized scientific research with their exceptional capabilities and transformed various fields.
-Among their practical applications, LLMs have been playing a crucial role in mitigating threats to human life, infrastructure, and the environment.
-Despite growing research in disaster LLMs, there remains a lack of systematic review and in-depth analysis of LLMs for natural disaster management.
-To address the gap, this paper presents a comprehensive survey of existing LLMs in natural disaster management, along with a taxonomy that categorizes existing works based on disaster phases and application scenarios.
-By collecting public datasets and identifying key challenges and opportunities, this study aims to guide the professional community in developing advanced LLMs for disaster management to enhance resilience against natural disasters.
+Large Language Models (LLMs) have demonstrated remarkable capabilities across various domains, including their emerging role in mitigating threats to human life, infrastructure, and the environment during natural disasters.
+Despite increasing research on disaster-focused LLMs, there remains a lack of systematic reviews and in-depth analyses of their applications in natural disaster management.
+To address this gap, this paper presents a comprehensive survey of LLMs in disaster response, introducing a taxonomy that categorizes existing works based on disaster phases and application scenarios.
+By compiling public datasets and identifying key challenges and opportunities, this study aims to provide valuable insights for the research community and practitioners in developing advanced LLM-driven solutions to enhance resilience against natural disasters.
 Harnessing Large Language Models for Disaster Management: A Survey
-Zhenyu Lei   Yushun Dong  Weiyu Li Rong Ding   Qi Wang   Jundong Li University of Virginia, Florida State University, Northeastern University {vjd5zr, jundong}@virginia.edu, yd24f@fsu.edu {weiy.li, ding.ro, q.wang}@northeastern.edu
+Zhenyu Lei   Yushun Dongthanks: Corresponding authors  Weiyu Li Rong Ding   Qi Wang11footnotemark: 1   Jundong Li11footnotemark: 1 University of Virginia, Florida State University, Northeastern University {vjd5zr, jundong}@virginia.edu, yd24f@fsu.edu {weiy.li, ding.ro, q.wang}@northeastern.edu
 1 Introduction
-Natural disasters have become increasingly frequent and severe, posing unprecedented threats to human life, infrastructure, and the environment Manyena (2006); Yu et al.
+Natural disasters are becoming increasingly frequent and severe, posing unprecedented threats to human life, infrastructure, and the environment Manyena (2006); Yu et al.
 (2018); Chaudhary and Piracha (2021).
-The 2010 Haiti earthquake, for instance, led to more than 200,000 fatalities and widespread infrastructure devastation DesRoches et al.
+The 2010 Haiti earthquake, for instance, resulted in over 200,000 fatalities and widespread infrastructure devastation DesRoches et al.
 (2011).
-Similarly, the 2020 Australian bushfires have resulted in the deaths of at least 33 people and an estimated one billion animals Deb et al.
+Similarly, the 2020 Australian bushfires caused the deaths of at least 33 people and an estimated loss of one billion animals Deb et al.
 (2020).
-The profound impacts of such catastrophic events have highlighted the urgent need for effective disaster management strategies.
-Recently, large language models (LLMs) have revolutionized the scientific research area due to their unparalleled abilities in contextual understanding, logical generation, and complex problem-solving with textual data and data from various modalities Zhang et al.
-(2024b, a).
-These capabilities make LLMs well-suited for natural disaster management, with models capable of analyzing vast disaster real-time data, facilitating dynamic communication with affected communities, and supporting critical decision-making Otal et al.
+The profound impact of such catastrophic events underscores the urgent need for effective disaster management strategies.
+Recently, large language models (LLMs) have transformed research and technological innovation with their exceptional capabilities in contextual understanding, logical reasoning, and complex problem-solving across multiple modalities Zhang et al.
+(2024b, a); Wei and Liu (2025).
+These capabilities position LLMs as powerful tools for natural disaster management, enabling them to analyze vast real-time disaster data, facilitate dynamic communication with affected communities, and support critical decision-making Otal et al.
 (2024).
-Despite the potential of LLMs in disaster management, a systematic review is still missing.
-This absence hinders researchers and practitioners from identify best practices, uncovering research gaps, and optimizing the deployment of LLMs for effectively addressing disaster-related challenges.
-To address this gap, this paper systematically surveys the applications of LLMs based on three types of architectures across the four disaster phases: mitigation, preparedness, response, and recovery.
-We then propose a novel taxonomy that integrates application scenarios, specific tasks, and the architectures of models addressing these tasks.
-Furthermore, we summarize publicly available datasets, identify key challenges, and investigate avenues for improving the effectiveness, efficiency, and trustworthiness of LLMs.
-This review aims to inspire and guide AI researchers, policymakers, and practitioners towards developing LLM-aided disaster management frameworks.
-Our contributions are listed below:
-  • We systematically introduce the current explorations handling disaster management with LLMs across four disaster phases.
-  • We propose a taxonomy integrating application scenarios, tasks, and model architectures, offering practical and technical insights.
-  • We collect key resources such as public datasets and highlight the current challenges and future research opportunities.
+Despite their potential, a systematic review of LLMs in disaster management remains absent, limiting researchers and practitioners in identifying best practices, addressing research gaps, and optimizing LLM deployment for disaster-related challenges.
+To bridge this gap, this paper presents a comprehensive survey of LLM applications in disaster management, categorizing them across three model architectures and the four key disaster phases: mitigation, preparedness, response, and recovery.
+We introduce a novel taxonomy that integrates application scenarios, specific tasks, and model architectures tailored to disaster-related challenges.
+Additionally, we summarize publicly available datasets, identify key challenges, and explore avenues for enhancing the effectiveness, efficiency, and trustworthiness of LLMs in disaster response.
+This review aims to inspire and guide AI researchers, policymakers, and practitioners toward developing LLM-driven disaster management frameworks.
+Our key contributions are as follows:
+  • Systematical Review: We provide the first systematical review of explorations of LLMs applications in disaster management across four key disaster phases.
+  • Novel Taxonomy: We propose a taxonomy integrating application scenarios, specific tasks, and model architectures, providing both practical and technical insights into this survey.
+  • Resource Compilation: We compile essential resources (e.g., datasets), and highlight key challenges and future research directions to advance LLM-driven disaster management.
+2 Background
+Disaster management is a multidisciplinary field that integrates resources, expertise, and strategies to mitigate the impact of increasingly severe disasters.
+Its primary goal is to minimize immediate damage while fostering long-term resilience and adaptive recovery.
+Disaster management comprises four interconnected phases Sun et al.
+(2020):
+  • Mitigation involves identifying risks and vulnerabilities while implementing proactive measures to prevent disasters.
+  • Preparedness includes developing comprehensive plans and public education initiatives to enhance readiness for potential disasters.
+  • Response identifies and addresses immediate needs during a disaster, including emergency rescue operations and resource distribution.
+  • Recovery involves rebuilding affected areas, addressing both physical and social impacts to facilitate a return to normalcy.
+In general, LLMs have the potential to serve as general-purpose foundations for developing specialized AI tools that enhance various aspects of disaster management.
+Here, we categorize LLM architectures into three main types: (1) encoder-based LLM (e.g., BERT Devlin (2018)), which excel in contextual understanding; (2) (encoder-)decoder LLM (e.g., GPT Brown (2020)), which are optimized for sequential prediction; and (3) multimodal LLMs, which integrate multiple modalities to enhance information processing Tiong et al.
+(2022); Madichetty et al.
+(2021) In disaster management, common downstream tasks include classification (e.g., damage classification), estimation (e.g., severity estimation), extraction (e.g., knowledge extraction), and generation (e.g., report generation).
+To tailor LLMs for these tasks, techniques such as fine-tuning and prompting are commonly employed.
+3 LLM For Disaster Management
+Foundation models can be utilized across the four disaster management phases: mitigation, preparedness, response, and recovery.
+Within each phase, existing works are categorized based on application scenarios, specific tasks, and model architectures.
+Figure 1 presents an overview of our taxonomy, with detailed summaries provided in Appendix A.
+3.1 Disaster Mitigation
+Assessing vulnerabilities is a crucial component of disaster mitigation, where LLMs have demonstrated promising potential.
+This process involves identifying and analyzing infrastructure and communities at risk, enabling proactive measures to reduce disaster impact.
+Vulnerability Classification.
+A system named Infrastructure Ombudsman has leveraged supervised learning with encoder-based LLMs and zero-shot prompt learning with (encoder-)decoder LLMs to detect and classify concerns about potential infrastructure failures from social media data Chowdhury et al.
+(2024).
+This approach enables decision-makers to effectively prioritize resources and address critical issues in a timely manner.
+Answer Generation.
+Beyond infrastructure vulnerability assessment, (encoder-)decoder LLMs can help address community vulnerability-related queries by retrieving and leveraging the Social Vulnerability Index (SVI) Martelo and Wang (2024).
+3.2 Disaster Preparedness
+In the long term, LLMs can play a pivotal role in disaster preparedness through (1) enhancing public awareness by disseminating accurate and accessible information, and (2) supporting disaster forecasting with advanced data analysis.
+Building on these forecasts, LLMs can aid decision-makers in issuing (3) timely disaster warnings, improving short-term preparedness.
+Furthermore, LLMs can support well-structured (4) evacuation planning, ensuring the safe relocation of individuals and assets.
 Figure 1: Taxonomy of applications of LLMs in disaster management.
-This survey categorizes the utilization of LLMs across four disaster phases, highlighting specific applications where three types of LLMs (Encoder-based, Decoder-based, and Multimodal LLM) perform tasks such as classification, estimation, extraction, and generation.
-The chart in the bottom-right corner presents the distribution of surveyed papers across each phase.
-2 Background
-Disaster management is a multidisciplinary field focused on coordinating resources, expertise, and strategies to address the challenges posed by increasingly severe disasters, minimizing immediate damage while fostering long-term resilience and adaptive recovery.
-Disaster management consists of four interconnected phases Sun et al.
-(2020):
-  • Mitigation. This phase involves identifying risks and vulnerabilities and implementing proactive measures for disaster prevention.
-  • Preparedness. This phase includes comprehensive planning and public education to ensure readiness for upcoming disasters.
-  • Response. This phase identifies and addresses immediate needs during a disaster, including rescue and resource distribution.
-  • Recovery. This phase involves rebuilding affected areas, addressing both physical and social impacts to facilitate a return to normalcy.
-In general, LLMs are able to serve as a general-purpose foundation for building specialized AI tools in disaster management.
-Here, we categorize LLM architectures into: (1) encoder-based LLM (e.g., BERT Devlin (2018)) primarily for extracting textual features for downstream tasks; (2) decoder-based LLM (e.g., GPT Brown (2020)) for text generation from given textual queries; (3) multimodal LLMs that integrate multiple modalities via alignment Tiong et al.
-(2022) or feature fusion Madichetty et al.
+This survey categorizes the utilization of LLMs across four disaster phases, highlighting specific applications where tasks such as classification, estimation, extraction, and generation are performed by three types of LLMs (Encoder-based, (Encoder-)Decoder, and Multimodal LLM).
+The chart in the bottom-right corner presents the distribution of papers across each phase.
+3.2.1 Public Awareness Enhancement
+Enhancing public awareness of disasters is crucial, particularly by providing insights and knowledge derived from past disaster experiences.
+Knowledge Extraction.
+Encoder-based LLMs have been fine-tuned to extract disaster-related knowledge from news articles and social media Fu et al.
+(2024), as well as from extensive disaster literature Zhang and Wang (2023), using Named Entity Recognition (NER).
+To improve the logical coherence of extracted entities, Ma et al.
+propose BERT-BiGRU-CRF for NER, enabling the construction of disaster knowledge graphs Ma et al.
 (2023).
-For disaster management, downstream tasks usually include classification (e.g., disaster detection), estimation (e.g., severity estimation), extraction (e.g., knowledge extraction), and generation (e.g., report generation).
-Techniques such as fine-tuning, prompting, and additional adapters are commonly employed to adapt LLMs for specific downstream applications.
-3 LLM For Disaster Management
-Foundation models can be applied to four phases of disaster management, namely mitigation, preparedness, response, and recovery.
-For each phase, works are organized by application scenarios, tasks, and model architectures.
-We present the taxonomy in Figure 1 and summaries in Appendix A.
-3.1 Disaster Mitigation
-Assessing vulnerabilities is a critical aspect of disaster mitigation, where LLMs have shown promising potential.
-This process typically involves identifying and analyzing infrastructure or communities that are susceptible to disasters.
-Vulnerability Classification.
-Infrastructure Ombudsman has compared the supervised classification with encoder-based LLMs and zero-shot prompt learning with decoder-based LLMs in detecting specific types of concerns about potential infrastructure failures from social media data Chowdhury et al.
-(2024), facilitating decision-makers in effectively prioritizing resources.
-Answer Generation.
-Apart from the infrastructure vulnerability, decoder-based LLMs can assist in answering community vulnerability-related queries based on retrieval of social vulnerability index (SVI) Martelo and Wang (2024).
-3.2 Disaster Preparedness
-In the long term, LLMs can play a pivotal role in disaster preparation via (1) public awareness enhancement and consistent (2) disaster forecast.
-Building on the forecasts, LLMs can support decision-makers in issuing (3) disaster warnings to the public, which can enhance preparedness in the short term.
-Furthermore, LLMs can assist in well-structured (4) evacuation planning, ensuring safe relocation of individuals and property.
-3.2.1 Public Awareness Enhancement
-It is critical to enhance public awareness regarding disasters by providing insights and knowledge extracted from previous disaster experiences.
-Knowledge Extraction.
-Encoder-based LLMs are particularly valuable due to their exceptional knowledge extraction capabilities.
-BERT has been fine-tuned to extract disaster-related knowledge from news and social media Fu et al.
-(2024), and extensive literature Zhang and Wang (2023).
-To enhance the logical coherence of extracted entities, Ma et al.
-propose BERT-BiGRU-CRF for name entity and relation recognition (NER) to construct disaster knowledge graphs Ma et al.
-(2023).
-With enhanced reasoning ability, decoder-based LLMs have been utilized, where ChatGLM2 is fine-tuned to extract knowledge triplets from documents to construct a knowledge graph Wu et al.
+In addition, (encoder-)decoder LLMs have been fine-tuned with instructional learning to extract knowledge triplets from documents for knowledge graph construction Wu et al.
 (2024).
 Answer Generation.
-The derived knowledge could be incorporated in decoder-based LLMs’ prompts, which could facilitate answering wildfire Hostetter et al.
-(2024) and flood risk-related questions Martelo and Wang (2024), and offering personalized mitigation guidance Li et al.
+The extracted disaster knowledge could be incorporated in (encoder-)decoder LLMs’ prompts, facilitating disaster-related question answering Hostetter et al.
+(2024); Martelo and Wang (2024); Li et al.
 (2023).
-In addition, techniques such as retrieval-augmented generation (RAG) have been utilized for knowledge incorporation, with knowledge retrieved from flood and geographic information system (GIS) knowledge graph Zhu et al.
+Additionally, techniques such as retrieval-augmented generation (RAG) have been employed to further improve knowledge integration Zhu et al.
 (2024).
-3.2.2 Disaster Forecast
-Effective disaster preparedness also hinges on accurate and robust forecasting and risk level prediction.
+3.2.2 Disaster Prediction
+Effective disaster preparedness also relies on accurate and reliable disaster prediction.
 Occurrence Classification.
-With their extensive internal knowledge, encoder-based LLMs have proven to be powerful tools for classification-based disaster forecasting.
-For instance, Indra et al.
-streamline BERT, GRU, and CNN to anticipate disaster Indra and Duraipandian (2023), improving their ability to capture temporal dependencies and local patterns in large-scale social media text.
-However, textual data is limited due to its subjective and imprecise nature, which leads to multimodal LLMs integrating various data modalities.
-Zeng et al.
-combine historical flood data and BERT embeddings extracted from geographical description, then utilize XGBoost to predict the disaster risk Zeng and Bertsimas (2023).
-Debata et al.
-further extract features from statistical data with k-nearest neighbors (KNN) for better incorporation.
-Further advancements incorporate additional satellite imagery features derived from ResNet for more informative disaster prediction Liu and Zhong (2023).
-To facilitate disaster risk prediction with explicit external knowledge, decoder-based LLMs have been utilized with RAG to retrieve historical flood data from disaster knowledge base to derive risk level and action recommendation Wang et al.
+Encoder-based LLMs have been widely employed for disaster prediction.
+For instance, BERT has been integrated with GRU and CNN to predict disasters Indra and Duraipandian (2023).
+However, textual data alone is often limited due to its subjective and imprecise nature, prompting the adoption of multimodal LLMs that incorporate multiple data modalities.
+For instance, Zeng et al.
+combine historical flood data with geographical descriptions of specific locations to assess disaster risk Zeng and Bertsimas (2023).
+Additionally, satellite imagery has been leveraged to provide visual context, enhancing predictive accuracy Liu and Zhong (2023).
+To further improve disaster prediction with explicit external knowledge, (encoder-)decoder LLMs have been integrated with RAG to retrieve historical flood data, aiding in risk assessment and action recommendation Wang et al.
 (2024).
 3.2.3 Disaster Warning
-Once a disaster has been foreseen, timely warnings and suggestions are crucial to public safety.
+Once a disaster is anticipated, timely warnings are essential for ensuring public safety.
 Warning Generation.
-Decoder-based LLMs have been a valuable tool in generating clear warning messages from rule-based alerts derived from streaming data Chandra et al.
-(2024), significantly enhancing the responsiveness of warning systems.
-In addition, LLMs have been utilized to retrieve disaster alerts from National Weather Service’s API to provide detailed, real-time information about upcoming disaster Martelo and Wang (2024).
-Image Generation In addition to textual warnings, visual warnings can convey vivid descriptions, reaching a broader audience.
-To achieve this, multimodal LLMs enhanced by diffusion-based text-to-image generative models can produce vivid visual representations of impending disasters Lubin et al.
-(2024), improving the clarity of alerts.
+(Encoder-)decoder LLMs have proven valuable in generating warning messages based on rule-based alerts derived from streaming data Chandra et al.
+(2024), significantly improving the responsiveness of warning systems.
+Additionally, RAG has enhanced LLMs by enabling the retrieval of disaster alerts from official APIs, providing real-time information on impending disasters Martelo and Wang (2024).
+Image Generation.
+In addition to textual warnings, visual warnings can provide more vivid and intuitive descriptions, effectively reaching a broader audience.
+To achieve this, multimodal LLMs enhanced by diffusion-based text-to-image generative models can generate detailed visual representations of impending disasters Lubin et al.
+(2024), enhancing the clarity and impact of disaster alerts.
 3.2.4 Evacuation Planning
-Plan Generation
-To ensure the safety of individuals and properties, structured evacuation plans have been developed.
-Leveraging the exceptional reasoning capabilities, decoder-based LLMs have been prompted to plan escapes and provide evacuation suggestions Hostetter et al.
-(2024), thus enhancing public safety against incoming disasters.
+Plan Generation.
+To safeguard individuals and property from impending disasters, (encoder-)decoder LLMs have been prompted to generate escape plans and provide evacuation recommendations Hostetter et al. (2024).
 3.3 Disaster Response
-With accurate and real-time (1) disaster identification and (2) situation assessment, decision makers can gain essential knowledge to form a response basis.
-In addition, LLM can facilitate the (3) disaster information coordination, thus boosting a collaborative disaster response among stakeholders.
-Consequently, decision makers can execute effective actions empowered by LLMs, such as (4) disaster rescue, and offering (5) disaster issue consultation.
+With accurate and real-time (1) disaster identification and (2) situation assessment, decision-makers can acquire critical insights to establish a solid foundation for response efforts.
+Additionally, LLMs can facilitate (3) disaster information coordination, enhancing collaboration among stakeholders for more effective disaster response.
+As a result, decision-makers can leverage LLMs to execute key actions, including (4) disaster rescue operations and (5) disaster-related consultations.
 3.3.1 Disaster Identification
-Effective disaster response begins with accurate and real-time identification to enable efficient interventions Said et al.
+Effective disaster response begins with accurate and real-time identification, enabling efficient interventions Said et al.
 (2019); Weber et al.
-(2020), with social media serving as a valuable avenue by providing real-time updates from affected individuals Anderson (2016); Trono et al.
 (2020).
-Relevance Classification. Leveraging their advanced capabilities, LLMs have become prominent disaster detection tools, particularly by performing binary classification for disaster-related posts.
-Encoder-based LLMs, with a trainable adapter appended, are widely utilized in this application via fine-tuning Ningsih and Hadiana (2021); Singh et al.
+Social media serves as a valuable resource in this process, offering real-time updates from affected individuals Anderson (2016); Trono et al.
 (2015).
-Beyond fine-tuning existing LLMs, researchers have developed novel techniques tailored to specific disaster contexts.
-CrisisTransformers has pre-trained transformers initialized randomly or with RoBERTa weights on a large dataset with millions of crisis-related tweets before fine-tuning for specific downstream tasks Lamsal et al.
-(2024a), thus enhancing extensive disaster knowledge incorporation.
-Additionally, Paul et al.
-tackle the challenge of limited labeled data by incorporating active learning, which fine-tunes transformer-based representations to prioritize the labeling of the most informative samples Paul et al.
+Relevance Classification with Encoder-based LLMs.
+Classifying social media posts to identify disaster-related content is a crucial step in disaster detection, where LLMs have proven to be highly effective.
+Encoder-based LLMs augmented with trainable adapters are commonly employed for this task through fine-tuning on annotated disaster corpora Ningsih and Hadiana (2021); Singh et al.
+(2022); Lamsal et al.
+(2024a).
+Recognizing the diverse sources of disaster data, ensemble methods combine predictions from multiple LLMs to leverage their complementary strengths in processing varied linguistic patterns Mukhtiar et al.
 (2023).
-More innovations focus on framework design, where ensemble models have been employed to integrate the strengths of various LLMs Mukhtiar et al.
+Pure LLM-based approaches may struggle to capture fine-grained structural features in disaster-related posts.
+To address this, hybrid architectures integrate CNNs to capture local n-gram patterns Franceschini et al.
+(2024); Song and Huang (2021); Meghatria et al.
+(2024), attention-based BiLSTMs to model sequential dependencies Huang et al.
+(2022), and graph neural networks (GNNs) to represent semantic word relationships Manthena (2023); Ghosh et al.
 (2022).
-Furthermore, additive modifications also enhance LLMs.
-For instance, several studies append CNN and att-BiLSTM to capture temporal dependencies and local patterns Manthena (2023).
-In addition, graph neural networks (GNN) can enhance word relation modeling.
-For instance, Danday et al.
-construct a word relationship graph from DistilBERT embeddings, which are sent to GNN for classification Danday and Murthy (2022).
-For better modality fusion, GNom integrates graph features derived from multilingual word graphs and BERT text embeddings with a novel cross-attention mechanism Ghosh et al.
+To tackle the challenge of limited labeled training data, active learning has been employed to automatically label informative samples Paul et al.
 (2023).
-In addition, advanced decoder-based LLMs such as Llama-2 and GPT-4 demonstrate exceptional contextual understanding ability, being prompted for classification Taghian Dinani et al. (2023).
-Image data also offer valuable insights, where multimodal LLMs—combining text and other images—has emerged as a promising approach, enhancing the accuracy of disaster identification.
-Several approaches combine text features from encoder-based LLMs and image features by aggregation Kamoji et al.
+Relevance Classification with Encoder-based LLMs.
+Furthermore, (encoder-)decoder LLMs such as GPT-4 have demonstrated strong performance in relevance classification using prompt learning techniques Taghian Dinani et al.
+(2023).
+Relevance Classification with Multimodal LLMs.
+Image data also provide valuable insights for disaster analysis and can be integrated to enhance classification using multimodal LLMs.
+This integration can be achieved through simple aggregation Kamoji et al.
 (2023); Madichetty et al.
+(2021) or attention-based mechanisms Shetty et al.
 (2024).
-In addition, Koshy et al.
-append a BiLSTM module after RoBERTa for better sequential modeling Koshy and Elango (2023).
-To improve modality alignment, Shetty et al.
-utilize the attention mechanism to fuse text and image features Shetty et al.
-(2024), while Zhou et al.
-adopt a Cycle-GAN with mixed fusion strategy to address the multimodal heterogeneity challenges Zhou et al.
+To address challenges arising from multimodal heterogeneity, Zhou et al.
+employ a Cycle-GAN combined with a mixed fusion strategy Zhou et al.
 (2023b).
-Beyond modality alignment, aligning features across various disaster events is critical in label-scarce scenarios.
-Jang et al.
-introduce a semi-supervised minimax entropy domain adaptation framework unifying multimodal features from various disaster events, enabling generalization across diverse contexts Wang and Wang (2022).
-To combine strengths of diverse LLMs and visual models, the ensemble is further utilized Hanif et al.
+Beyond multimodal heterogeneity, research also tackles other challenges in multimodal learning.
+These include addressing label scarcity through semi-supervised minimax entropy domain adaptation frameworks Wang and Wang (2022) and enhancing model performance by leveraging the complementary strengths of diverse LLMs and visual models using ensemble methods Hanif et al.
 (2023).
-In addition to social media data, other real-time data sources, such as satellite imagery and news articles, can also provide valuable insights.
-For instance, Jang et al.
-utilize ResNet to process satellite images and BERT to process descriptions of the disasters from news Jang et al.
+Beyond social media, data from sources such as satellite imagery and news articles can further enhance disaster analysis Jang et al.
 (2024).
 3.3.2 Disaster Situation Assessment
-After identifying a disaster, evaluating its severity and spread is crucial for response strategies.
-This information is often embedded in textual data, such as social media posts reporting situation Saroj and Pal (2020), and news reports offering precise descriptions Houston et al.
-(2012); Sood et al.
+After disaster identification, assessing its severity and spread is essential for formulating effective response strategies.
+Situation Classification.
+encoder-based LLMs have been fine-tuned to for binary classification to identify situational posts Madichetty and Sridevi (2021).
+Raj et al.
+employ BERT and NER to extract disaster-related locations, using location counts as an indicator of disaster severity Raj et al.
 (2023).
-Situation Classification.
-Madichetty et al.
-fine-tune encoder-based LLMs to identify situational posts Madichetty and Sridevi (2021), combining the RoBERTa and SVM representations for binary classification.
-Raj et al.
-utilize BERT to filter disaster-related tweets, followed by NER to extract locations, with location counts indicating disaster severity Raj et al.
-(2023).
-In addition, visual data from the web and satellites is also integrated and processed by multimodal LLMs to enhance the disaster assessment.
-Kanth et al.
-utilize BERT and CNN to filter disaster-related tweets and classify the disaster severity based on the images associated with informative tweets Kanth et al.
+Additionally, multimodal LLMs integrate visual data to further enhance disaster situational assessment Kanth et al.
 (2022).
 Severity Estimation.
-Classification only offers a coarse understanding, while severity estimation provides precise quantitative insights.
-Mousavi et al.
-utilize chain-of-thoughts (CoT) prompting on decoder-based LLMs (e.g.
-Gemini) to estimate the earthquake intensity as Modified Mercalli Intensity (MMI) values Mousavi et al.
+While classification provides only a coarse understanding, severity estimation offers precise quantitative insights.
+(encoder-)decoder LLMs enhanced with chain-of-thought (CoT) reasoning have been used to estimate earthquake intensity, expressed as Modified Mercalli Intensity (MMI) Mousavi et al.
 (2024).
-In addition, multimodal LLMs can be utilized to perform image-related estimation.
-For instance, FloodDepth-GPT uses prompts to guide GPT-4 to estimate the floodwater depth of the floodwater Akinboyewa et al.
-(2024) from flood images and descriptions.
+In addition, multimodal LLMs use rich image data for more accurate estimations.
+For example, FloodDepth-GPT employs prompt-based guidance with GPT-4 to estimate floodwater depth from flood images.
 Description Generation.
-Beyond categorical and statistical descriptions, multi-modal LLMs can generate more comprehensible textual situational descriptions.
-For instance, GPT-4V is prompted to translate situational images to descriptions Hu and Rahnemoonfar (2024), while Wolf et al.
-prompts GPT-4 to generate assessment descriptions Wolf et al.
-(2023) with disaster image data.
+Beyond categorical and statistical descriptions, multimodal LLMs can generate more comprehensible textual situational reports from disaster images Hu and Rahnemoonfar (2024); Wolf et al.
+(2023).
 3.3.3 Disaster Information Coordination
-Coordinating disaster-related information is essential for an organized, collaborative response Comfort et al.
+Coordinating disaster-related information is crucial for ensuring an organized and collaborative response Comfort et al.
 (2004); Bharosa et al.
 (2010).
-Social media plays a pivotal role, with individuals sharing posts on warnings, needs, and other types of information Lindsay (2011); Imran et al.
+Social media plays a pivotal role in this process, as individuals actively share posts containing warnings, urgent needs, and other critical information Lindsay (2011); Imran et al.
 (2015).
 Usefulness Estimation.
-To enhance the accessibility of useful information, Yamamoto et al.
-leverage encoder-base LLMs (e.g.
-BERT and RoBERTa) to filter informative tweets by calculating usefulness ratings Yamamoto et al.
+To improve the accessibility of valuable information, encoder-base LLMs are utilized to filter informative tweets by computing usefulness ratings Yamamoto et al.
 (2022).
+However, this approach requires a predefined threshold to determine the relevance of a tweet.
 Relevance Classification.
-However, it requires a pre-defined threshold to determine tweet relevance.
-Several studies fine-tune encoder-based LLMs for binary relevance classification, which can also be regarded as disaster identification.
-Furthermore, BERT is applied to classify posts by relevance with multiple relevance levels Blomeier et al.
+Several studies fine-tune encoder-based LLMs for binary relevance classification, as discussed in Section 3.3.1.
+Additionally, LLMs have been applied to multi-level relevance classification to further refine disaster-related information filtering Blomeier et al.
 (2024).
 Information Classification.
-To facilitate information dissemination, several studies fine-tuned encoder-based LLMs to classify posts by information types, including actionable types such as "important for managers" Sharma et al.
+To facilitate information dissemination, several studies have fine-tuned encoder-based LLMs to classify posts based on different information types, including actionable types such as "important for managers" Sharma et al.
 (2021); humanitarian types such as "Injured people" Yuan et al.
-(2022); and disaster types Liu et al.
+(2022); and disaster-specific types Liu et al.
 (2021).
-Several studies introduce advanced techniques to improve performance, with some focusing on augmenting data.
-Boros et al.
-augment tweets with hashtags and entities, leveraging RoBERTa for classification and tweet prioritizing Boros et al.
+When fine-tuning data is limited, augmentation strategies such as manual hashtag annotation Boros et al.
+(2022) and self-training with soft labeling Li et al.
+(2021) are employed to enhance classification performance Lei et al.
 (2025).
-Addressing data label scarcity, Li et al.
-introduce self-training with BerTweet Nguyen et al.
-(2020) soft-labeling data for training Li et al.
+Pure LLM-based methods may have limitations, as discussed in Section 3.3.1.
+In contrast, hybrid architectures enhance performance by integrating CNNs and BiLSTMs to improve local pattern comprehension Zou et al.
+(2024) and employing Graph Attention Networks (GATs) to capture correlations between tweet embeddings and information types Zahera et al.
 (2021).
-Novel frameworks have been introduced to improve performance.
-For instance, Zou et al.
-utilize a BERT-Bi-LSTM-CNN model to classify tweets into damage types Zou et al.
-(2024), with better abilities in capturing temporal dependencies and local patterns.
-In addition, Zahera et al.
-utilize a graph attention network (GAT) to capture correlations between tweet embeddings and information types Zahera et al.
+Additionally, FF-BERT leverages an ensemble of BERT and CNN to combine model strengths for improved classification Wilkho et al.
 (2024).
-Furthermore, FF-BERT combines model strengths with an ensemble of BERT and CNN Wilkho et al.
+Other studies enhance the application of LLMs in disaster information classification by extracting rationales—evidence that supports classification decisions Nguyen and Rudra (2022b, 2023).
+RACLC Nguyen and Rudra (2022a) employs a two-stage framework, utilizing contrastive learning to refine rationale extraction and improve classification performance.
+(Encoder-)decoder LLMs have also been employed for disaster type and humanitarian classification through instruction tuning Otal and Canbaz (2024); Yin et al.
+(2024), as well as zero-shot and few-shot prompting Dinani et al.
 (2024).
-Other works enhance LLMs by introducing inductive bias.
-Nguyen et al.
-leverage BERTweet to extract rationales (classification supporting evidence) from tweets for humanitarian classification Nguyen and Rudra (2022b) and integrate them with attention mechanisms for better fusion Nguyen and Rudra (2023).
-RACLC Nguyen and Rudra (2022a) further adopts a two-stage framework, using contrastive learning to refine rationale extraction and facilitate classification.
-Furthermore, advanced decoder-based LLMs have been utilized, through instruction-tuning Otal and Canbaz (2024); Yin et al.
-(2024), zero-shot, and few-shot prompting Dinani et al.
-(2024) for disaster type and humanitarian classification.
-Visual data from social media is further incorporated by multimodal LLMs, where several studies have combined features from BERT and visual models via aggregation Zhang et al.
-(2022); Yu and Wang (2024).
-Better fusion strategies have been proposed, where Abavisan et al.
-propose a cross-attention mechanism to fuse DenseNet image features with BERT text embeddings for humanitarian classification Abavisani et al.
+Multimodal LLMs can integrate rich data from social media to enhance classification by leveraging multiple modalities Lei et al.
 (2022).
-Zhou et al.
-advance this approach by utilizing dual transformers to fuse ALBERT text features and S-CBAM-VGG image features Zhou et al.
+This integration can be achieved through simple feature aggregation Zhang et al.
+(2022); Yu and Wang (2024) or more advanced fusion techniques, such as cross-attention mechanisms Abavisani et al.
+(2020) and dual transformer architectures Zhou et al.
 (2023a).
-In addition, Basit et al.
-classify posts into humanitarian or structural types using the "OR" set of text and image classification results Basit et al.
+Additionally, Basit et al.
+classify posts into humanitarian or structural categories only when the text and image classification outputs align; otherwise, the posts are uninformative Basit et al.
 (2023).
 Need Classification.
-Social media also allows individuals to express needs, inspiring LLMs for need identification.
-Encoder-based LLMs have been utilized to detect disaster needs Yang et al.
-(2024) and rescue request Toraman et al.
+Social media enables individuals to express urgent needs during disasters.
+Encoder-based LLMs have been employed to detect disaster-related needs Yang et al.
+(2024); Vitiugin and Purohit (2024) and rescue requests Toraman et al.
 (2023).
-Furthermore, MulTMR employs multiple teacher models with task knowledge to fine-tune BERT and RoBERTa for ranking and detecting actionable help requests Vitiugin and Purohit (2024).
-Responders also share resources on social media.
-Lamsal et al.
-have adopted LLMs to match needs with resources through cosine similarity-based retrieval methods where both offer and request posts are embedded by XLM-RoBERTa Conneau (2019), optimizing resource allocation Lamsal et al.
-(2024b).
+Responders also use social media to share available resources.
+Encoder-based LLMs have been employed to match needs with resources using cosine similarity-based retrieval methods, where both offer and request posts are embedded using XLM-RoBERTa Conneau (2019), optimizing resource allocation.
 Location Extraction.
-Additionally, efficient processing of posts enhances information dissemination, especially through location extraction.
-Several studies fine-tune encoder-based LLMs for location reference recognition (LRR), classifying tokens into categories such as "Inside Locations" (ILOC), and "Other Tokens"(O) Mehmood et al.
+Additionally, various post-processing techniques enhance information dissemination, particularly through location extraction.
+Several studies fine-tune encoder-based LLMs for location reference recognition (LRR), classifying tokens into categories such as "Inside Locations" (ILOC) and "Other Tokens" (O) Mehmood et al.
 (2024); Suwaileh et al.
-(2022).
-Koshy et al.
-further improve LRR by adding attention and Bi-LSTM layersKoshy and Elango (2024), while others use a BERT-BiLSTM-CRF model to improve the logic consistency of extracted locations Ma et al.
+(2022); Koshy and Elango (2024).
+LRR can be further improved by integrating a conditional random field (CRF) model, which enhances the logical consistency of extracted locations Ma et al.
 (2022); Zhang et al.
 (2021).
-External knowledge bases further aid location extraction.
+Furthermore, external knowledge corpora can support location extraction.
 For instance, Caillaut et al.
-utilize cosine similarity to match post entities with knowledge base Caillaut et al.
-(2024), improving performance.
-Decoder-based LLMs are also widely applied, where Yu et al.
-prompt Llama-3 to extract location-relevant words Yu and Wang (2024).
-External knowledge has been utilized for more accurate location extraction.
-Hu et al.
-leverage GPT-4 for detailed location descriptions, incorporating geo-knowledge within the prompts Hu et al.
+use cosine similarity to match post entities with a knowledge base, ensuring the authenticity of extracted locations Caillaut et al.
 (2024).
-In addition, GeoChatGPT augments LLMs with geographic data and Object Character Recognition (OCR)-based object descriptions to predict disaster locations Firmansyah et al.
+(Encoder-)decoder LLMs are widely used for extracting location-relevant information through prompt learning Yu and Wang (2024).
+To enhance accuracy, external knowledge has been incorporated into prompts, including geo-knowledge Hu et al.
+(2023) and Object Character Recognition-based object descriptions Firmansyah et al.
 (2024).
 Summary Extraction.
-Furthermore, extracting and summarizing critical posts provides a macro understanding during disasters.
-Nguyen et al utilize integer linear programming (ILP) with encoder-based LLMs to identify critical posts and rationales for summaries Nguyen and Rudra (2022a); Nguyen et al.
+Furthermore, summarizing disaster-related posts provides a macro-level understanding during crises.
+Several studies focus on identifying critical and informative posts for summarization by integrating advanced techniques into encoder-based LLMs, such as integer linear programming (ILP) Nguyen and Rudra (2022a); Nguyen et al.
+(2022) and Rapid Automatic Keyword Extraction (RAKE) Garg et al.
 (2024).
-Garg et al.
-calculate salience scores using BERT embeddings and word2vec key-phrase embeddings from rapid automatic keyword extraction (RAKE)Garg et al.
-(2024).
-Decoder-based LLMs go further by generating summaries from retrieved tweets.
-For instance, Vitiugin et al.
-rank important tweets utilizing LSTM and then use a T5 model to generate the summaries based on top retrieved tweets Vitiugin and Castillo (2022).
+Summary Generation.
+(Encoder-)decoder LLMs extend summarization capabilities by generating summaries from retrieved text.
+For example, Vitiugin et al.
+rank key tweets using an LSTM model and apply a T5 model to generate summaries based on the top-ranked tweets Vitiugin and Castillo (2022).
+Crisis2Sum performs query-focused summarization through a multi-step process, including query-informed document retrieval, reranking, fact extraction, clustering, fusion into event nuggets, and final selection for summarization Seeberger and Riedhammer (2024a).
+Additionally, agent-based approaches can enhance summary quality by leveraging multiple LLMs for document retrieval, reranking, and instruction-following summarization Seeberger and Riedhammer (2024b).
 Report Generation.
-Organizations also use decoder-based LLMs for structured disaster reporting, especially through zero-shot search-based generation Colverd et al.
-(2023).
-In addition, Seeberger et al.
-utilize multiple LLMs each for document retrieval, reranking, and instruction-following summarization Seeberger and Riedhammer (2024b), while chain-of-thought (CoT) reasoning is further utilized Pereira et al.
-(2023).
-Building on these methods, Crisis2Sum further organizes query-related information into topic clusters summarized by Mixtral Jiang et al.
-(2024), and produces a final report by combining key nuggets from each cluster Seeberger and Riedhammer (2024a).
+(Encoder-)decoder LLMs have been employed for disaster report generation, utilizing techniques such as RAG to extract relevant web data Colverd et al.
+(2023) and Chain-of-Thought reasoning to enhance the coherence and accuracy of generated reports.
 3.3.4 Disaster Rescuing
-Grounded in a comprehensive disaster situation understanding, disaster rescuing aims to save lives and property through timely, coordinated actions.
+Grounded in a comprehensive understanding of the disaster situation, disaster rescue focuses on saving lives and protecting property through timely and coordinated actions.
 Plan Generation.
-Effective rescue operations typically require comprehensive rescue plans.
-Grounded with rich disaster-related knowledge, decoder-based LLMs have been prompted to generate actionable response plans that provide essential guidance Goecks and Waytowich (2023).
+Effective rescue operations require well-structured rescue plans.
+(Encoder-)decoder LLMs are prompted to generate actionable response plans, offering essential guidance for disaster response Goecks and Waytowich (2023).
 Code Generation.
-Once a plan is established, decoder-based LLMs can assist in execution performed by organizations and rescue teams, such as assisting in guiding robotic systems by translating verbal inputs into actionable operational commands using RAG Panagopoulos et al.
+Once a plan is established, (encoder-)decoder LLMs can support its execution by assisting organizations and rescue teams.
+For instance, they can facilitate robotic system guidance during rescue operations by translating verbal inputs into actionable operational commands using RAG Panagopoulos et al.
 (2024).
@@
 During disasters, affected individuals and organizations often seek reliable guidance.
-Disaster issue consultation offers advice, safety updates, and expert guidance, thus helping them access resources, evaluate options, and make decisions Jiang (2024).
+Disaster issue consultation provides advice, safety updates, and expert recommendations, helping them access resources, evaluate options, and make informed decisions Jiang (2024).
 Answer Generation.
-Decoder-based LLMs have been prompted to generate answers for frequently asked questions and offer guidance Rawat (2024); Chen and Fang (2024).
-To mitigate hallucination, LLMs integrate RAG with verified disaster-related documents.
-For instance, WildfireGPT retrieves wildfire literature and data to augment the prompts Xie et al.
+(Encoder-)decoder LLMs are employed to generate answers for frequently asked questions and provide disaster-related guidance Rawat (2024); Chen and Fang (2024).
+To mitigate hallucination, RAG is integrated with verified disaster-related documents.
+For example, WildfireGPT retrieves wildfire-related literature and data to enhance prompts Xie et al.
 (2024).
 Chen et al.
-guide LLM reasoning over KG by a prompt chain, adding structured knowledge Chen et al.
+introduce a prompt chain to guide LLM reasoning over a disaster knowledge graph, incorporating structured knowledge Chen et al.
 (2024).
-Unlike traditional RAG approaches that rely on prompt learning, Xia et al.
-integrate fine-tuning with RAG to enhance LLMs by combining implicit and external knowledge.
-Xia et al.
+Unlike traditional RAG approaches without training, Xia et al.
+combine fine-tuning for implicit knowledge updates with RAG for explicit knowledge, further improving response quality Xia et al.
 (2024).
-Additionally, multi-modal LLMs can integrate textual and visual data.
-For instance, several VQA models (e.g., Plug-and-Play VQA) Tiong et al.
-(2022) have been prompted for zero-shot visual question answering for disaster scenarios Sun et al.
+Additionally, multi-modal LLMs can integrate textual and visual data to enhance disaster response.
+For example, several visual question answering (VQA) models, such as Plug-and-Play VQA Tiong et al.
+(2022), have been prompted for zero-shot VQA in disaster scenarios Sun et al.
 (2023).
-To interpret complex user queries, ADI further introduces sequential modular tools including VLMs, object detection models, and semantic segmentation models Liu et al.
+To handle complex user queries, ADI introduces sequential modular tools, incorporating vision-language models (VLMs), object detection models, and semantic segmentation models Liu et al.
 (2024).
-In addition, FloodLense integrates ChatGPT and diffusion models to highlight disaster-affected areas from images to improve flood-related geographical question answering Kumbam and Vejre (2024).
+Furthermore, FloodLense combines ChatGPT with diffusion models to highlight disaster-affected areas in images, enhancing flood-related geographical question answering Kumbam and Vejre (2024).
 3.4 Disaster Recovery
-LLMs can perform accurate (1) disaster impact assessment, a critical step for disaster recovery.
-With a comprehensive impact understanding, LLMs can assist decision makers in (2) recovery plans generation.
-In addition, disaster responders have utilized LLMs for (3) recovery process tracking consistently, ensuring the recovery effectiveness.
+LLMs can play a crucial role in (1) disaster impact assessment, a vital step in the recovery process.
+By providing a comprehensive understanding of disaster impacts, LLMs can assist decision-makers in (2) generating recovery plans tailored to specific needs.
+Additionally, disaster responders have leveraged LLMs for (3) continuous recovery process tracking, ensuring effectiveness and progress throughout the recovery phase.
 3.4.1 Disaster Impact Assessment
-Accurate assessment of the damage degree across physical and social dimensions helps prioritize recovery efforts effectively.
-Physically, disasters disrupt infrastructure, damage property, and threaten human lives, requiring robust frameworks for the identification and estimation of damage and loss.
+Accurately assessing the extent of damage across both physical and social dimensions is essential for prioritizing recovery efforts effectively.
 Damage Classification.
-Various encoder-based LLMs have been employed to identify and categorize disaster damage.
-For instance, Malik et al.
-utilize BERT to identify the damage-related posts and categorize them as "human damage" or "infrastructure damage" Malik et al.
+From the physical dimension, encoder-based LLMs have been employed to identify and categorize disaster-related damage (e.g., human/infrastructure damage Malik et al.
+(2024), water/power supply damage Chen and Lim (2021)) Additionally, Zou et al.
+propose a BERT-BiLSTM-Sit-CNN framework, improving textual understanding for damage-related post identification and damage-type classification Zou et al.
 (2024).
-Chen et al.
-further categorize damage into four types such as "Water or Power Supply" with BERT Chen and Lim (2021).
-In addition, Zou et al.
-propose a BERT-BiLSTM-Sit-CNN framework for damage-related posts identification and damage-type classification with improved textual understanding Zou et al.
-(2024).
-Furthermore, several studies utilize LLMs to measure damage severity, where Jeba et al.
-leverage BERT for damage impact degree classification on social media posts and news  Jeba et al.
+Beyond type classification, LLMs have been utilized to assess damage severity.
+For instance, Jeba et al.
+employ BERT to classify damage impact severity in social media posts and news articles Jeba et al.
 (2024).
 Damage Estimation.
-Damage severity could be better modeled quantitatively by estimation with a fine-grained understanding.
+Damage severity can be more effectively quantified through fine-grained estimation.
 Chen et al.
-calculate the damage severity scores based on the similarities between post tokens and several seed words’ embeddings, where both embeddings are derived from encoder-based LLMs Chen and Lim (2021).
+compute damage severity scores by measuring the similarity between post tokens and predefined seed words’ embeddings, both of which are derived from encoder-based LLMs Chen and Lim (2021).
 Answer Generation.
-In addition, decoder-based LLMs can answer specific assessment questions.
-For instance, Ziaullah et al.
-prompt LLMs with RAG to obtain the operation status of critical infrastructure facilities using social media data Ziaullah et al.
+In addition, (encoder-)decoder LLMs can answer specific assessment questions.
+Ziaullah et al.
+employ RAG-enhanced LLMs to retrieve operational status updates of critical infrastructure facilities from social media data Ziaullah et al.
 (2024).
-Multimodal LLMs further incorporate remote sensing data, with Estevao et al.
-prompting GPT-4o to generate damage assessment using building images Estêvão (2024).
-To improve modality alignment, SAM-VQA employs a supervised attention-based VLM to integrate image and question features for VQA tasks Sarkar et al.
+Multimodal LLMs further incorporate remote sensing data for enhanced assessment.
+Estevao et al.
+prompt GPT-4o to generate damage assessments based on building images Estêvão (2024).
+To improve modality alignment, SAM-VQA employs a supervised attention-based vision-language model (VLM) to integrate image and question features for visual question answering (VQA) tasks Sarkar et al.
 (2023).
-Additionally, axillary tasks have been utilized to improve the VQA performance, where DATWEP dynamically balances the importance of segmentation and VQA tasks, adjusting class weights during training Alsan and Arsan (2023).
-Statistic Extraction. Furthermore, decoder-based LLMs are prompted to extract fatalities Hou and Xu (2022) from social media via few-shot learning, thus offering timely insights into human loss.
+Additionally, auxiliary tasks have been leveraged to enhance VQA performance.
+For instance, DATWEP dynamically balances the significance of segmentation and VQA tasks by adjusting class weights during training Alsan and Arsan (2023).
+Statistic Extraction. (Encoder-)decoder LLMs have also used few-shot learning to extract fatality information from social media Hou and Xu (2022), offering timely insights into human loss.
 Sentiment Classification.
-From the social dimension, disasters influence public sentiment and behavior, where encoder-based LLMs Han et al.
+From the social dimension, disasters can influence public sentiment, where encoder-based LLMs Han et al.
 (2024a); Berbère et al.
@@
 In addition, Li et al.
-utilize decoder-based LLM (e.g.
-GPT 3.5) to classify the posts into five emotional types such as "panic" and "sadness" with zero-shot prompting Li et al.
-(2025), thus aiding responders in understanding and addressing the disaster’s emotional impact.
+employ (encoder-)decoder LLM (e.g., GPT 3.5) to classify posts into five emotional types, such as "panic" and "sadness", using zero-shot prompting Li et al.
+(2025).
+This approach helps responders better understand and address the emotional impact of disasters.
 3.4.2 Recovery Plan Generation
-Based on impact assessment, a recovery plan is developed to rebuild infrastructure, restore services, and enhance resilience Hallegatte et al. (2018).
+Based on impact assessment, a recovery plan is formulated to rebuild infrastructure, restore services, and strengthen resilience Hallegatte et al. (2018).
 Plan Generation.
-Decoder-based LLMs have been utilized in certain recovery scenarios by generating recovery and reconstruction plans.
-For instance, ChatGPT has been prompted to build disaster recovery plans for business restoration White and Liptak (2024); Lakhera (2024).
+(Encoder-)decoder LLMs have been applied in certain recovery scenarios to generate recovery and reconstruction plans.
+For example, ChatGPT has been prompted to develop disaster recovery strategies for business restoration White and Liptak (2024); Lakhera (2024).
+Figure 2: A sample of dataset for disaster relevance classification from CrisisLexT6 Olteanu et al. (2014).
 3.4.3 Recovery Process Tracking
-Continuous tracking of the recovery process ensures that progress aligns with the planned timeline and goals and decision-makers can adapt recovery strategies to meet evolving needs.
+Continuous tracking of the recovery process ensures that progress remains aligned with the planned timeline, allowing decision-makers to adapt recovery strategies to evolving needs.
 Sentiment Classification.
-Encoder-based LLM (e.g.
-BERTweet) has been utilized to assess public sentiment throughout the post-disaster recovery period CONTRERAS et al.
-, thus enabling responders to tailor their recovery efforts to address emotional needs of affected population effectively.
+Encoder-based LLMs (e.g., BERTweet) have been employed to assess public sentiment throughout the recovery period CONTRERAS et al.
+, enabling responders to tailor recovery efforts to effectively address the emotional needs of affected populations.
+Figure 3: Publication counts utilizing existing LLMs and developing new models across the four phases of disaster management.
 4 Datasets
-Numerous disaster-related datasets have been utilized to evaluate LLMs in disaster management, and we present a comprehensive list of publicly available datasets in Appendix B.
-Classification datasets are primarily designed with textual inputs from platforms like Twitter and news outlets, categorizing data by informativeness (relevance) Olteanu et al.
-(2014), humanitarian types Imran et al.
+Multiple disaster-related datasets have been employed to evaluate LLMs in disaster management. A comprehensive list of publicly available datasets is provided in Appendix B.
+Classification datasets primarily consist of textual inputs from platforms such as Twitter and news outlets, categorizing data based on informativeness (relevance) Olteanu et al.
+(2014) (illustrated in Figure 2), humanitarian types Imran et al.
 (2016), damage levels Alam et al.
-(2021b), and other attributes.
-Some also include visual data, such as satellite imagery or social media images Alam et al.
+(2021b), and other relevant attributes.
+Some datasets also incorporate visual data, including satellite imagery and social media images Alam et al.
 (2018).
-Models are evaluated with metrics like accuracy and F1 score.
-Estimation datasets include textual and visual inputs and provide quantitative labels such as flood depths Akinboyewa et al. (2024), evaluated using metrics such as Mean Absolute Error (MAE).
-Generation datasets are also extensively used, primarily in two categories: question answering and summarization.
-Question-answering datasets provide disaster-related questions paired with crowd-source annotated answers Rawat (2024).
-Multimodal question-answering datasets, which incorporate disaster-related images as contextual information, are widely utilized as well Sun et al.
+Model performance is typically evaluated using metrics such as accuracy and F1 score.
+Estimation datasets usually provide quantitative labels such as flood depths Akinboyewa et al. (2024). Metrics like Mean Absolute Error (MAE) are used for evaluation.
+Generation datasets are also extensively used and primarily fall into two categories: question answering and summarization.
+Question-answering datasets provide disaster-related questions paired with crowdsourced annotated answers Rawat (2024).
+Additionally, multimodal question-answering datasets, which incorporate disaster-related images as contextual information, are widely utilized Sun et al.
 (2023).
-For summarization task, large sets of documents are used as inputs, with reference summaries curated by domain experts McCreadie and Buntain (2023).
-Both types of generation tasks are evaluated using metrics like ROUGE and BLEU.
-Extraction datasets identify and label specific elements within sentence, such as keywords Nguyen and Rudra (2022a) and locations Suwaileh et al.
-(2022), labeling tokens as “outside,” “start,” or “end” to specify their extraction status.
-These datasets are primarily used for token-level classification tasks, and is measured with classification metrics.
+For the summarization task, large collections of documents serve as inputs, with reference summaries curated by domain experts McCreadie and Buntain (2023).
+Both question-answering and summarization tasks are evaluated using metrics such as BLEU.
+Extraction datasets identify and label specific elements within a sentence, such as keywords Nguyen and Rudra (2022a) and locations Suwaileh et al.
+(2022).
+Tokens are labeled as "outside," "start," or "end" to indicate their extraction status.
+These datasets are primarily used for token-level classification tasks and are evaluated using classification metrics.
 5 Challenges and Opportunities
-Large language models (LLMs) show promise in disaster management but face key limitations.
-Applications are disproportionately focused on disaster response, as illustrated in Figure 1, highlighting the need to expand their use across all disaster phases.
-Additionally, there is a gap between researchers focusing on novel architectures and practitioners needing disaster-specific solutions, necessitating LLMs tailored to various challenges.
-Dataset Scarcity.
-Current datasets are heavily skewed toward classification tasks, leaving other critical areas underrepresented.
-Creating high-quality datasets requires significant effort and resources, making the process both labor- and data-intensive.
-Innovative solutions like synthetic data generation Kalluri et al.
-(2024) are promising to expand coverage across disaster-related scenarios.
+Large Language Models (LLMs) hold great promise for disaster management but face several key limitations.
+Most studies deploy generic LLMs as universal solutions, overlooking domain-specific challenges and the need for tailored frameworks, as shown in Figure 3.
+Additionally, current applications are heavily concentrated on disaster response, leaving other phases underexplored, as illustrated in Figure 1.
+To fully harness the potential of LLMs in disaster management, researchers must address the disaster-specific challenges outlined below.
+Dataset Construction.
+Current datasets are heavily skewed toward classification tasks, leaving other areas underexplored Proma et al.
+(2022).
+Additionally, raw disaster data often contains uncertainty and bias Smith and Katz (2013), posing challenges in constructing reliable datasets.
+Innovative approaches, such as synthetic data generation Kalluri et al.
+(2024), offer a promising solution to enhance dataset coverage across diverse disaster scenarios.
 Efficient Deployment.
-large-scale LLMs face latency and computational challenges, hindering real-time decision-making in emergent disaster scenarios.
-Lightweight models provide a faster, resource-efficient alternative Saleem et al.
-(2024), but they often sacrifice accuracy and robustness.
-Advancing these models to balance efficiency with reliability is crucial for effective disaster management.
+Large-scale LLMs face efficiency challenges Ramesh Raja et al.
+(2024), limiting their viability for real-time decision-making in emergency disaster scenarios.
+While lightweight models offer a more efficient alternative Saleem et al.
+(2024), they often compromise robustness in disaster-related tasks.
+Developing models that balance efficiency and reliability is essential for effective disaster management.
 Robust Generation.
-Decoder-based LLMs may hallucinate, where outputs can be factually inaccurate, posing serious risks in disaster contexts, such as false evacuation routes, resource misallocation, and potential loss of lives.
-Integrating RAG with external knowledge bases Colverd et al.
-(2023), extensive domain-specific training Lamsal et al.
+(Encoder-)decoder LLMs are prone to hallucination, generating factually inaccurate outputs that pose serious risks in disaster contexts, such as false evacuation routes, resource misallocation, and potential loss of lives.
+To mitigate these risks, strategies such as integrating RAG with external knowledge bases Colverd et al.
+(2023), domain-specific training Lamsal et al.
 (2024a), and uncertainty estimation Xu et al.
-can potentially deal with hallucinated outputs.
+can help reduce hallucinated outputs and improve reliability.
+Unified Evaluation.
+Most generative benchmarks for disaster (e.g.
+report/summary generation) rely on reference sets produced by domain experts.
+While these curated answers provide high-quality supervision, the underlying annotation criteria such as what counts as a correct answer often differ from one dataset to another.
+Consequently, published results are difficult to compare directly, because each study is implicitly tied to its own expert standard.
+As a result, it’s important to build up a unified evaluation protocol to make more reliable comparisons.
 6 Conclusion
-This paper surveys LLMs in disaster management across four disaster phases, with a taxonomy integrating application scenarios, specific tasks, and the architectures of models addressing these tasks.
-With introducing publicly available datasets and identifying challenges, we inspire the collaborative efforts between AI researchers and decision-makers, in order to fully leveraging the power of LLMs for more resilient communities and proactive disaster management practices.
+This paper surveys the application of LLMs in disaster management across the four disaster phases, introducing a taxonomy that integrates application scenarios, specific tasks, and the architectures of models addressing these tasks.
+By presenting publicly available datasets and identifying key challenges, we aim to inspire collaborative efforts between AI researchers and decision-makers, ultimately enabling the full potential of LLMs to build more resilient communities and advance proactive disaster management practices.
+Acknowledgment
+This work is supported in part by the National Science Foundation (NSF) under grants IIS-2006844, IIS-2144209, IIS-2223769, CNS-2154962, BCS-2228534, BCS-2228533, CMMI-2411248, CMMI-2125326, and CMMI-2402438.
+All authors would like to thank the reviewers and chairs for their constructive feedback and suggestions.
 Limitations
-This work focuses exclusively on disaster management applications only where existing LLMs have been utilized, leaving out other potential scenarios that have not yet been explored by current LLM research.
-While these unexplored areas hold significant promise for future advancements, they are beyond the scope of this work due to space constraints.
-Additionally, we include only a subset of datasets employed in existing studies, prioritizing those that are easily accessible.
-Many datasets are either not open-sourced, have restrictive access policies, or lack assured quality, making them less viable for reproducibility and further research.
+Survey Scope.
+This work focuses exclusively on disaster management applications only where existing LLMs have been utilized, leaving out other potential scenarios (e.g.
+repair cost evaluation during the recovery phase) that have yet to be explored in current research.
+In addition, we focus mostly on the natural disaster instead of man-made disaster.
+While these unexplored areas hold significant promise for future advancements, they fall beyond the scope of this study due to space constraints.
+Categorization.
+This work categorizes papers based on their model architecture.
+However, it would be also beneficial to analyze existing papers from the other perspectives such as model size, inference efficiency, and model performance.
+Datasets.
+Additionally, we include only a subset of datasets used in existing studies, prioritizing those that are easily accessible.
+Many datasets either are not open-sourced, have restrictive access policies, or lack assured quality, making them less suitable for reproducibility and further research.
 Appendix A Summary of Papers
 A.1 Summary Table
-Table LABEL:tab:model provides a summary of the surveyed papers with their disaster phases, application scenarios, specific tasks, and architecture types.
+Table LABEL:tab:model summarizes the surveyed papers, detailing their disaster phases, application scenarios, specific tasks, and architecture types.
 A.2 Pipeline Illustration
-In this section, we present Figure 2 to illustrate the use of LLMs in disaster management.
-The figure outlines the major pipelines of three LLM architectures—encoder-based, decoder-based, and multimodal—applied across four task types in our survey: classification, extraction, estimation, and generation, offering critical insights into their mechanisms for disaster management.
+In this section, we present Figure 4, which illustrates the role of LLMs in disaster management.
+The figure outlines the major pipelines of three LLM architectures—encoder-based, (encoder-)decoder, and multimodal—applied across the four task types covered in this survey: classification, extraction, estimation, and generation.
+This visualization provides key insights into their mechanisms and applications in disaster management.
 A.3 Statistics
 To provide a comprehensive overview of the current state of LLMs in disaster management, we present statistics from the surveyed papers, highlighting a significant gap between the NLP and disaster management communities.
-This gap underscores the pressing need for better interdisciplinary collaboration to bridge the two fields and realize the potential of LLMs in addressing disaster-related challenges.
-Figure 3 depicts the counts of publications leveraging existing LLMs and developing new frameworks, which indicates that most studies are heavily application-focused, relying on fine-tuning or prompting existing LLMs for disaster management tasks.
-In addition, while some efforts have yielded valuable insights, the majority of these studies are predominantly concentrated on the response phase of disaster management, with limited exploration across other disaster management scenarios.
-Figure 4 illustrates the publication distribution across academic venues, revealing that relatively few disaster management papers are published in NLP or AI-specific conferences or journals.
-These findings reflect limited engagement from the LLM research community in this domain, highlighting a need to increase awareness and focus within the field.
+This gap underscores the urgent need for stronger interdisciplinary collaboration to bridge these fields and fully harness the potential of LLMs in addressing disaster-related challenges.
+Figure 3 illustrates the number of publications leveraging existing LLMs versus those developing new frameworks, revealing that most studies are heavily application-focused.
+The majority rely on fine-tuning or prompting existing LLMs for disaster management tasks, rather than designing novel architectures.
+While some efforts have provided valuable insights, most research remains concentrated on the response phase, with limited exploration across other critical disaster management scenarios.
+Figure 5 illustrates the distribution of publications across academic venues, revealing that relatively few disaster management papers appear in NLP- or AI-specific conferences and journals.
+This trend reflects limited engagement from the LLM research community in this domain, underscoring the need to increase awareness and foster greater collaboration within the field.
 Appendix B Datasets
-Table LABEL:tab:dataset provides a summary of existing publicly available datasets.
-For classification tasks, we exclude datasets limited to a single disaster type that are already incorporated into comprehensive benchmarks such as CrisisBench Alam et al.
+Table LABEL:tab:dataset summarizes existing publicly available datasets.
+For classification tasks, we exclude datasets that focus on a single disaster type if they are already incorporated into comprehensive benchmarks such as CrisisBench Alam et al.
 (2021b).
@@
   • CrisisNLP Imran et al.
-(2016): This dataset is a large-scale dataset including classes from humanitarian disaster responses and classes related to health emergencies.
+(2016): This dataset is a large-scale dataset that includes classes from humanitarian disaster responses and classes related to health emergencies.
 It is collected from 19 different disaster events that happened between 2013 and 2015.
   • SWDM2013 Imran et al.
-(2013): This dataset is utilized for relevance classification consisting of tweets from two events: (i) the Joplin collection contains tweets from the tornado that struck Joplin, Missouri on May 22, 2011; (ii) The Sandy collection contains tweets collected from Hurricane Sandy that hit Northeastern US on Oct 29, 2012.
+(2013): This dataset is utilized for relevance classification that consists of tweets from two events: (i) the Joplin collection contains tweets from the tornado that struck Joplin, Missouri on May 22, 2011; (ii) The Sandy collection contains tweets collected from Hurricane Sandy that struck the Northeastern US on Oct 29, 2012.
   • ISCRAM2013 Imran et al. (2013): This dataset consists of tweets collected from the same events as in SWDM2013, containing both relevance and humanitarian categories.
@@
   • Eyewitness Messages Zahra et al.
-(2020): This dataset is proposed to identity disaster eyewitness related tweets and classifies them into three types: direct eyewitnesses, indirect eyewitnesses, and vulnerable eyewitnesses where people anticipate a disaster and were present in the region for which disaster warnings were issued.
-It is comprised of $14k$ data from earthquakes, hurricanes, and wildfires.
+(2020): This dataset is designed to identify disaster eyewitness-related tweets and classify them into three categories: direct eyewitnesses, indirect eyewitnesses, and vulnerable eyewitnesses—individuals who anticipate a disaster and are present in regions where disaster warnings have been issued.
+It comprises 14,000 tweets collected from earthquakes, hurricanes, and wildfires.
   • TREC Incident Streams McCreadie et al.
@@
 It is used for relevance and humanitarian classification.
-Figure 3: Counts of publications leveraging existing LLMs and developing new models across four phases.
+Figure 5: Publication distribution across various academic venues, with a focus on model design on the left and application-based research on the right.
 B.2 Estimation Datasets
-  • Did You Feel It (DYFI) Atkinson and Wald (2007): This dataset contains ground shaking intensity and geographic distribution information, collected from post-earthquake reports (through $750k$ online questionnaire responses) from individuals who experienced the event.
+  • Did You Feel It (DYFI) Atkinson and Wald (2007): This dataset includes ground shaking intensity and geographic distribution information, collected from post-earthquake reports through 750,000 online questionnaire responses from individuals who experienced the event.
   • FloodDepth Akinboyewa et al.
-(2024): This dataset contains 150 flood photos collected online, which is utilized to estimate floodwater depth based on different reference objects, including stop signs, vehicles, and humans.
+(2024): This dataset consists of 150 flood photos collected online, used to estimate floodwater depth based on various reference objects, including stop signs, vehicles, and humans.
   • Behavioral Facilitation (BF) Yamamoto et al.
-(2022): This datasets, collected after the 2018 Hokkaido earthquake, contains data labeled with usefulness ratings with respect to behavioral facilitation information.
-Figure 4: Publication distribution across various academic venues, with a focus on model design on the left and application-based research on the right.
+(2022): This dataset, collected after the 2018 Hokkaido earthquake, includes data labeled with usefulness ratings based on behavioral facilitation information.
 B.3 Extraction Datasets
@@
 It is utilized to extract information such as the time and location of disasters.
-  • Ma et al. (2023): This dataset is used for entity and relation extraction, containing $5,560$ annotated data, $12,980$ entities, and $6,895$ relations, from report during geological hazards.
+  • Ma et al.
+(2023): This dataset is designed for entity and relation extraction, comprising 5,560 annotated instances, 12,980 entities, and 6,895 relations derived from reports on geological hazards.
   • DisasterMM Andreadis et al.
-(2022): This dataset is collected from Twitter by searching for flood-related keywords.
-It comprises two subsets: RCTP, containing 6,672 tweets for relevance classification; LETT, consisting of 4,992 tweets used for location extraction, where words are labeled with “B-LOC” for the first word of a sequence that refers to a location, “I-LOC” for the subsequent word of a sequence that refers to a location, and “O” for any non-location word.
+(2022): This dataset was collected from Twitter by searching for flood-related keywords.
+It consists of two subsets: RCTP, which includes 6,672 tweets for relevance classification, and LETT, which contains 4,992 tweets used for location extraction.
+In the LETT subset, words are annotated with "B-LOC" for the first word in a sequence referring to a location, "I-LOC" for subsequent words within the same location sequence, and "O" for words that do not correspond to a location.
   • Suwaileh et al.
-(2022): This dataset contain $22k$ crisis-related tweets from several disasters including floods, earthquakes, and hurricane.
-It is labeled with location-related tags such as "inLOC" and "outLOC".
+(2022): This dataset contains 22,000 crisis-related tweets from various disasters, including floods, earthquakes, and hurricanes.
+It is annotated with location-related tags such as "inLOC" and "outLOC."
   • Re´SoCIO Caillaut et al.
@@
   • Nguyen and Rudra (2022a): This dataset contains tweet data with annotated rationales from 4 subsets of CrisisNLP.
-It is utilized for rationale extraction, and the extracted rationales can assist in disaster classification.
+It is used for rationale extraction, and the extracted rationales can assist in disaster classification.
 B.4 Generation Datasets
-  • Vitiugin and Castillo (2022): This dataset is utilized to generate summaries for various disaster events, where the ground truth is the official report of each event.
-  • CrisisFACTS McCreadie and Buntain (2023): This dataset is a multiple-stream dataset with data related to 8 crisis events from various platforms.
-It is utilized to consume daily multi-platform streams and produce summaries for a given information need, such as "Have airports closed?".
+  • Vitiugin and Castillo (2022): This dataset is used to generate summaries of various disaster events, with the official report of each event serving as the ground truth.
+  • CrisisFACTS McCreadie and Buntain (2023): This dataset is a multi-stream collection comprising data from eight crisis events gathered across various platforms.
+It is designed to process daily multi-platform streams and generate summaries based on specific information needs, such as "Have airports closed?"
   • DisasterQA Rawat (2024): This dataset includes disaster-related multiple choice questions from 7 different sources, examples could be "What causes a tsunami?".
   • FFD-IQA Sun et al.
-(2023): This dataset includes 2,058 images and 22,422 question-meta ground truth pairs related to the safety of individuals trapped in disaster sites and the availability of emergency services.
-In total, three types of questions are included: free-form questions, multiple choice questions, and yes-no questions.
+(2023): This dataset comprises 2,058 images and 22,422 question-meta ground truth pairs related to the safety of individuals trapped in disaster sites and the availability of emergency services.
+It includes three types of questions: free-form, multiple-choice, and yes-no questions.
   • FloodNet Rahnemoonfar et al.
-(2021): The dataset contains 4,500 question-image pairs collected after Hurricane Harvey.
-Questions are related to the building, road, and the entire image, and are divided into 4 groups: "Simple Counting", "Complex Counting", "yes/no", and "Condition Recognition".
+(2021): This dataset consists of 4,500 question-image pairs collected after Hurricane Harvey.
+The questions pertain to buildings, roads, and entire scenes, categorized into four groups: "Simple Counting," "Complex Counting," "Yes/No," and "Condition Recognition."
 Table 1: Summary of LLMs in disaster management with their disaster phases, application scenarios, specific tasks, and architecture types.
@@
 Singh et al. (2022) Response Disaster Identification Relevance Classification Encoder No
-Ranade et al. (2021) Response Disaster Identification Relevance Classification Encoder No
 Powers et al. (2023) Response Disaster Identification Relevance Classification Encoder No
-Jain et al. (2019) Response Disaster Identification Relevance Classification Encoder No
 Duraisamy and Natarajan (2024) Response Disaster Identification Relevance Classification Encoder No
@@
 Nguyen and Rudra (2022b) Response Disaster Information Coordination Information Classification Encoder Yes
+Zou et al. (2023) Response Disaster Information Coordination Information Classification Encoder Yes
 Nguyen and Rudra (2023) Response Disaster Information Coordination Information Classification Encoder Yes
 Nguyen and Rudra (2022a) Response Disaster Information Coordination Information Classification Encoder Yes
+Dar et al. (2025) Response Disaster Information Coordination Information Classification Encoder Yes
 Otal and Canbaz (2024) Response Disaster Information Coordination Information Classification Decoder No